### v1

# Train Once, Deploy Anywhere: Realize Data-Efficient Dynamic Object Manipulation

Abstract
Realizing generalizable dynamic object manipulation is important for enhancing manufacturing efficiency, as it eliminates specialized engineering for various scenarios. To this end, imitation learning emerges as a promising paradigm, leveraging expert demonstrations to teach a policy manipulation skills. Although the generalization of an imitation learning policy can be improved by increasing demonstrations, demonstration collection is labor-intensive. To address this problem, this paper investigates whether strong generalization in dynamic object manipulation is achievable with only a few demonstrations. Specifically, we develop an entropy-based theoretical framework to quantify the optimization of imitation learning. Based on this framework, we propose a system named Generalizable Entropy-based Manipulation (GEM). Extensive experiments in simulated and real tasks demonstrate that GEM can generalize across diverse environment backgrounds, robot embodiments, motion dynamics, and object geometries. Notably, GEM has been deployed in a real canteen for tableware collection. Without any in-scene demonstration, it achieves a success rate of over 97% across more than 10,000 operations.
Index Terms:
Dynamic object manipulation, imitation learning, information entropy, manufacturing efficiency.
I Introduction
Robotic manipulation of dynamic objects has broad applications in industrial manufacturing, particularly in conveyor-based assembly lines where robots handle transported products [1, 2]. Compared with the extensively studied static object manipulation problem [3, 4], dynamic object manipulation poses more challenges due to its demand for more time-efficient motion planning and stricter operational precision. Addressing these challenges is important for enhancing the efficiency and flexibility of industrial automation systems, thereby enabling more economical manufacturing.
Currently, the methods employed in real-world production lines are primarily based on classic motion planning approaches [5]. These methods require specialized designs for various scenarios, as well as complex system modeling and parameter tuning, which could be both time-consuming and labor-intensive [6]. To handle this problem, much attention has been drawn to exploring how to build an intelligent manipulation system capable of generalizing across diverse scenarios. Recent manipulation literature increasingly turns to the imitation learning paradigm [7], where expert demonstrations are collected to teach a policy to control a robot arm. In imitation learning, the generalization of a policy can be improved by scaling up the volume of demonstrations [8]. Consequently, the research community has devoted significant efforts to constructing large-scale manipulation datasets [9, 10]. However, these datasets primarily focus on static scenarios, with few public datasets addressing dynamic object manipulation. Additionally, even in static object manipulation, the generalization of the policies trained using these large-scale datasets is still limited [11, 12], implying the volume of demonstration is still far from sufficient.
Therefore, although realizing generalizable dynamic object manipulation is of huge importance, collecting demonstrations is costly and there are few public data resources [13]. Given these facts, this work studies whether strong generalization in dynamic object manipulation can be achieved without relying on numerous demonstrations. To address this question, we delve into the optimization essence of imitation learning, framing it as a distribution approximation problem. Building upon this formulation, we propose an information entropy based theoretical framework to quantify the amount of data required to achieve generalizable manipulation. From this theoretical framework, we identify two key aspects: first, reducing the complexity of redundant information in observations, and second, decreasing the ambiguity in action prediction.
Figure 1: Primarily using demonstrations collected from a simulator, our method can generalize across diverse environment backgrounds, robot embodiments, motion dynamics, and object geometries. Our method has been reliably deployed in a real canteen to conduct tableware collection. Without relying on any demonstration collected in this canteen, our method achieves a success rate of over 97% in seven consecutive days of operation, performing more than 10,000 tableware collection operations.
According to these two key aspects, we develop a manipulation system named Generalizable Entropy-based Manipulation (GEM). GEM consists of two blocks, i.e., object-centric geometry perception and hybrid action control. In object-centric geometry perception, we devise a representation that emphasizes the geometric structures of objects while minimizing the impact of visual appearance, thereby reducing redundant details irrelevant to manipulation. Within hybrid action control, imitation learning is combined with visual servo control to manage the moving dynamics of objects. Additionally, an efficient memory modeling strategy is proposed in the imitation learning algorithm to alleviate action prediction ambiguity.
In experiments, we implement a single GEM to support four common skills in dynamic object manipulation, i.e., Pick, Put, Rotate, and Insert. Extensive experiments are conducted in both simulators and the real world to validate the effectiveness of GEM. The experiments suggest that GEM presents promising generalization across diverse environment backgrounds, robot embodiments, motion dynamics, object geometries, and in sim-to-real settings. Especially, as illustrated in Fig. 1, GEM has been successfully deployed in a real canteen to perform tableware collection. Without using demonstrations collected in this canteen, GEM achieves a 97.2% success rate over seven days of testing and more than 10,000 operations.
Summarily, the contributions of this work are as follows:
1) We develop an entropy-based framework to analyze the optimization of imitation learning, guiding the design of an efficient manipulation system with limited demonstrations.
2) We propose GEM, a system combining object-centric geometry perception and hybrid action control to achieve strong generalization in dynamic object manipulation.
3) We demonstrate the first real-world application of imitation learning in dynamic object manipulation, achieving over 97% success rate without in-scene demonstrations.
II Related Works
II-A Imitation Learning for Manipulation
Traditional robotic manipulation methods mostly rely on accurate modeling of the environment to plan action trajectories of robot arms [5]. However, manipulation often involves contact-rich interactions with objects, and accurately modeling these interactions is difficult due to the highly variable nature of contact forces [14]. Moreover, every manipulation task often requires a unique set of models and parameters, making these traditional methods difficult to generalize across different objects, environments, and robot embodiments. This necessitates substantial engineering efforts to adapt the system to new tasks or scenarios [15].
Given these limitations of traditional methods, there is a growing interest in leveraging imitation learning to address the challenges of robotic manipulation [7]. Bypassing the need for explicit system modeling and complex parameter tuning, imitation learning learns manipulation policies directly from expert demonstrations. Specifically, imitation learning trains a policy to predict expert actions based on the observations provided in demonstrations, enabling a robot to mimic expert behavior in similar scenarios. The policy is often implemented as a deep learning network [16], which can capture the complex patterns in demonstrations [17].
The two most commonly employed observation modalities in robotic manipulation are images and 3D points [5]. Correspondingly, existing imitation learning policies can be broadly categorized into image-based [18] and point-based approaches [19]. Images are easy to obtain, and image-based policies can readily leverage large-scale public data for training [20]. However, robotic manipulation mainly concerns the geometric structures of objects, while images primarily capture the appearance of scenes, making it challenging to predict appropriate manipulation actions based solely on visual appearance. In contrast, point clouds directly represent the geometric structures of scenes, and thus they are more suitable for regressing manipulation actions [21]. Unfortunately, point cloud data is relatively scarce compared with images, resulting in a lack of promising point cloud feature extractors [22]. This limitation has historically constrained the performance of point-based manipulation policies [23]. Thankfully, some recent works about large-scale point cloud pre-training have gradually addressed this obstacle [24, 25].
II-B Data Scaling in Manipulation
The rapid progress in fields like large language models [26] suggests that increasing the quantity and diversity of training data is important for enhancing the generalization of deep learning models. Drawing on this insight, the robotic manipulation community has invested significant effort into constructing large-scale demonstration datasets [9, 10]. Leveraging these datasets, numerous studies design algorithms for training end-to-end manipulation policies, with the expectation that these policies exhibit robust generalization across diverse scenarios [11, 27]. Furthermore, some works incorporate additional types of data like object detection datasets to further enrich the knowledge base of manipulation policies [12, 28]. Nevertheless, though much resource has been invested, the generalization capabilities of the obtained policies in unseen environments are still limited.
This fact reminds us to rethink whether the data scaling paradigm inspired by large language models is directly applicable to robotic manipulation. Rather than solely focusing on constructing larger manipulation datasets or training models with more parameters, we also need to pay attention to designing an efficient manipulation paradigm that achieves economical manipulation generalization. This shift in perspective emphasizes the importance of developing methods that can leverage limited data more efficiently [29].
II-C Sim-to-Real Generalization in Manipulation
Collecting high-quality demonstration data in the real world is labor-intensive. To address this issue, many researchers have explored generating demonstrations in simulated environments [30]. These efforts aim to develop policies trained using only simulated demonstrations yet can generalize to real-world applications. A key first step towards this goal is the development of realistic simulators. Through substantial engineering efforts, several promising simulators have been built, such as MuJuCo [31], Isaac Gym [32], and Isaac Sim [33]. Leveraging these simulators, prior works have used reinforcement learning [34] or predefined kinematic procedures [35] to generate demonstrations. However, reinforcement learning often struggles to discover successful manipulation trajectories due to its inefficient exploration [36]. Grasping objects via predefined kinematic procedures typically requires customized workflows for different scenarios, which is time-consuming. Additionally, current simulators have limited environmental diversity, restricting the variety of generated data.
Another line of works focuses on modeling the environment dynamics of manipulation tasks using classic mechanics and then employs optimization techniques to calculate actions that satisfy task constraints [14]. The observation and action pairs can then be utilized to train deep neural networks [37, 38]. This strategy effectively addresses the slow speed of optimization by distilling the action search process into a more efficient neural network [39]. Nevertheless, for complex manipulation tasks, the computation burden is too heavy to search for satisfactory actions, making it difficult to generate effective demonstrations. To tackle this problem, some works collect demonstrations in simulators through teleoperating simulated robots [40]. Although this method still requires significant human effort, it reduces hardware setup costs compared to real-world data collection.
III Entropy Reduction Theoretical Framework
This section develops a theoretical framework by analyzing the information entropy of imitation learning, providing logical support for developing a generalizable manipulation system with only a few demonstrations.
III-A Robotic Manipulation Problem Formulation
Robotic manipulation involves controlling a robot arm to interact with objects in the environment to achieve a specific task. The task can be defined as a sequence of actions that the robot needs to take to reach a desired goal. Therefore, the robotic manipulation problem can be conceptualized as a Markov Decision Process represented by a tuple $\mathcal{E}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{T},s_{0})$. In this tuple, $\mathcal{S}$ denotes the set including all potential states of the environment and robot. $\mathcal{A}$ represents the actions that can be performed by the robot. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the probability distribution governing how the current state transitions to the next state given the performed action. $\mathcal{T}$ represents the task target set and $s_{0}$ denotes the initial state.
In a specific task target $\tau\in\mathcal{T}$ of robotic manipulation, there usually exist numerous accurate action trajectories that can lead the robot from the initial state $s_{0}$ to the desired goal state $\iota$. We denote all the accurate trajectories as a set $\mathcal{C}$ and each trajectory is represented as a sequence of states and actions $\{s_{0},a_{0},\cdots,s_{T-1},a_{T-1},s_{T}\}$, where $s_{T}=\iota$.
To enable the robot to automatically complete the target task $\tau$ without human intervention, researchers devise algorithms to find an optimal policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the probability of reaching the target state $\iota$ following the transition probability distribution $\mathcal{P}$. This objective can be expressed as learning to approximate the joint state-action distribution induced by the successful trajectories in $\mathcal{C}$. Specifically, let $p_{\mathcal{C}}(s_{0:T},a_{0:T-1})$ denote the state-action sequence distribution in $\mathcal{C}$, the optimization objective is to minimize the Kullback-Leibler (KL) divergence between the policy predicted distribution $p_{\pi}(s_{0:T},a_{0:T-1})$ and $p_{\mathcal{C}}(s_{0:T},a_{0:T-1})$:
$\min\limits_{\pi}D_{KL}(p_{\mathcal{C}}(s_{0:T},a_{0:T-1})\ ||\ p_{\pi}(s_{0:T},a_{0:T-1})),$ (1)
where $p_{\pi}(s_{0:T},a_{0:T-1})$ is defined through the Markov chain:
$p_{\pi}(s_{0:T},a_{0:T-1})=p(s_{0})\prod\limits_{t=0}^{T-1}\pi(a_{t}|s_{t})P(s_{t+1}|s_{t},a_{t}).$ (2)
This formulation explicitly connects policy optimization with implicit distribution estimation of $\mathcal{C}$.
III-B Entropy Analysis of Imitation Learning
However, in robotic manipulation, a policy $\pi$ usually does not have access to the true state $s_{t}$ at timestamp $t$. Its environment observation is a mapping function of $s_{t}$, written as $o_{t}=f(s_{t})$. Compared with $s_{t}$, $o_{t}$ only partially reflects the environment information and may contain noise. The action to execute is sampled from $\pi$ following $a_{t}\sim\pi(o_{t},\tau)$. The policy $\pi$ is often modeled as a deep neural network [16].
Many recent manipulation works derive $\pi$ based on imitation learning [8], which is typically formulated as a supervised learning problem. Imitation learning requires access to a set of expert demonstrations as training data. Assume there are $N_{d}$ demonstration sequences, the expert demonstration set is denoted as $\mathcal{C}_{d}=\{o_{0:T}^{i},\tilde{a}_{0:T}^{i}\}_{i=1}^{N_{d}}$, where $o_{t}^{i}$ and $\tilde{a}_{t}^{i}$ are the observation and expert action at timestamp $t$ of the $i_{\rm th}$ demonstration. The core idea of imitation learning is employing $\mathcal{C}_{d}$ as a surrogate of $\mathcal{C}$. Thus, the objective of imitation learning can be expressed as maximizing the likelihood of expert demonstrations under the policy-induced distribution:
$\max\limits_{\pi}\mathbb{E}_{(o_{t}^{i},\tilde{a}_{t}^{i})\sim\mathcal{C}_{d}}[\log\pi(\tilde{a}_{t}^{i}|o_{t}^{i},\tau)].$ (3)
This objective can be understood as that $\pi$ is trained to capture the joint observation-action probability distribution in $\mathcal{C}_{d}$.
In practical implementations of imitation learning, the objective in Eq. (3) is often transformed into a behavior cloning objective, which is to explicitly minimize the difference between the predicted actions and the expert actions, written as:
$\pi^{*}=\arg\min\limits_{\pi}\sum\limits_{i=1}^{N_{d}}\sum\limits_{t=0}^{T}L(\pi(o_{t}^{i},\tau),\tilde{a}^{i}_{t}),$ (4)
where $L$ represents the difference measurement function, and $\pi^{*}$ is the obtained optimal policy. Inspired by the recent breakthroughs in research fields like large language models [26] and image generation [41], it is believed that the distribution of $\mathcal{C}$ can be precisely approximated by scaling up the volume of expert demonstration data $\mathcal{C}_{d}$ [11, 9].
Nevertheless, collecting diverse demonstrations is expensive in robotic manipulation [42]. Although huge efforts have been invested in building large-scale manipulation datasets [9, 10], the available quantity of demonstrations is far from sufficient compared with fields like large language models [43]. Currently, the policies trained on these datasets only serve as pre-trained network weights, still requiring task-specific fine-tuning for each downstream application [27, 42]. Thus, exploring strategies to reduce the demonstration data requirement of imitation learning is important.
To conduct this exploration, we first specify that the optimization objective of imitation learning is to approximate the joint observation-action distribution $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$, which describes how to achieve successful manipulation. For notational convenience, we represent the sets of observations and actions in $\mathcal{C}$ as $\mathcal{O}_{\mathcal{C}}$ and $\mathcal{A}_{\mathcal{C}}$. To explore how to reduce the demonstration demand while maintaining policy performance, we need a theoretical framework to quantify the data requirement for approximating $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$. In information theory, the entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}},\mathcal{A}_{\mathcal{C}})$ serves as a measure of the amount of information contained within a distribution, reflecting the inherent uncertainty or complexity of data [44]. A higher entropy indicates the distribution contains more information and requires more demonstrations to accurately approximate.
Furthermore, we decompose $\mathcal{H}(\mathcal{O}_{\mathcal{C}},\mathcal{A}_{\mathcal{C}})$ via the chain rule:
$\mathcal{H}(\mathcal{O}_{\mathcal{C}},\mathcal{A}_{\mathcal{C}})=\mathcal{H}(\mathcal{O}_{\mathcal{C}})+\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}}).$ (5)
This decomposition reveals that the information entropy of imitation learning consists of two parts, the observation complexity $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and the action prediction ambiguity $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$. By reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, the demonstration volume requirement can be alleviated.
In addition, we expect the manipulation performance not to compromise after reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$. To figure out how to realize this goal, it should be first noted that although there are numerous manipulation trajectories meeting constraints, $\pi$ only needs to predict one of them to complete the task. Therefore, when reducing the entropy, the probability of selecting a suitable action leading to successful manipulation should not be decreased. To ensure this, (i) When lessening the observation complexity $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, we only remove redundant observation information irrelevant to inferring accurate actions. (ii) During reducing the action prediction ambiguity $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, the policy should increase the likelihood of predicting a successful trajectory.
Figure 2: This figure illustrates the primary factors affecting the observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. The factors include environment, robot, sensor, and task instruction. These factors can be further divided.
III-C Observation Complexity Reduction
To alleviate the demonstration data demand, we first study how to reduce the observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, which measures the complexity of the observation set in $\mathcal{C}$. Specifically, we expect to reduce the redundant information in observations that is irrelevant to inferring actions, because removing such information does not hurt action prediction precision. Notably, which kinds of information are irrelevant to actions are not deterministic. They are influenced by how the manipulation system is designed. For instance, if a policy $\pi$ directly predicts the robot hand end-effector movement and then utilizes this movement to control joint rotations based on inverse kinematics, the joint angle observation is unnecessary for inferring actions when the end-effector position observation has been provided. Therefore, the key to reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is devising a manipulation system making most factors irrelevant to inferring actions. Then, these irrelevant factors can be excluded from observations, thereby reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
To more thoroughly analyze this problem, we need to know $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is determined by which factors. We illustrate the primary factors in Fig. 2. These main factors contributing to $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ can be categorized into four classes, i.e., environment information, robot proprioception, sensor configuration, and task instruction. Among them, the environment information includes the target object to manipulate, scene background, and ambient conditions (such as lighting brightness and air humidity). These pieces of information can be further divided, like the target object information can be classified into its physical geometry and visual appearance.
Differently, robot proprioception mainly describes the type of robot employed (like Franka Panda or UR5e), where the robot is located in the environment, and the action space (like end-effector control or joint control). Sensor configuration specifies what kinds of sensors are adopted (like RGB camera and depth sensor), how the sensors are deployed in the environment (the number of sensors and their poses), and the parameters of these sensors. A task instruction specifies the task target. It can be represented in different modalities, including text [11], image [40], voice [45], etc.
Summarily, the above four classes of factors play distinct roles in shaping the complexity of the observation set $\mathcal{O}_{\mathcal{C}}$. To reduce $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, we need to design a system making most factors irrelevant to predicting actions. We explain how such a specific system is devised in Section IV.
Figure 3: The sub-figure (a)-(c) illustrate the procedures that a robot hand presses a button. We can find that if the policy input does not include historical memory and moving velocity, the policy observations are the same between sub-figure (a) and (c), in which the robot hand is approaching the button and moving away from the button, respectively.
III-D Action Prediction Ambiguity Reduction
In this subsection, we discuss how to reduce $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ in Eq. (5), which represents the action prediction ambiguity given observations. This ambiguity is mainly caused by the multimodal nature of robotic manipulation, meaning there exist multiple valid action trajectories given the same observation [46]. For example, in the task of grasping a block, a robot arm can complete the task by approaching the block from either the left or the right side. Moreover, the ambiguity becomes more pronounced when the observation information is insufficient. For instance, Fig. 3 illustrates an example that a robot hand is pressing a button. In Fig. 3 (a)-(c), the robot hand is approaching the button, pressing the button, and departing from the button, respectively. If the policy input only includes the current visual observation and robot joint angles, the observations are the same between Fig. 3 (a) and (c), while the accurate end-effector moving directions are different.
Hence, larger action prediction ambiguity not only increases demonstration demand, but also raises the risk of making incorrect decisions. To handle this problem, more comprehensive information supporting accurate action prediction should be provided. According to the example in Fig. 3, incorporating historical memory into policy observation is effective, especially for complex and long-chained manipulation tasks. This is because historical memory enables the policy to know which manipulation steps have been completed.
However, previous works often find incorporating historical memory degrades success rates [8]. One key reason is that these works choose to model memory using historical action trajectories. As analyzed in VIP [40], when the attribute to predict (the future action trajectory in this case) and a part of the policy input (the historical action trajectory) fall in the same modality, the policy often chooses to overlook the input information in other modalities (like images). This is because it is easier to predict future actions based on historical actions than information in different modalities. A learnable network tends to choose the easiest way of regressing future actions while overlooking the exploitation of information in other modalities. This problem is termed as shortcut learning and has been well analyzed in previous literature [47].
An effective strategy for handling the shortcut learning problem is preventing using input in the same modality as output. In this case, it means avoiding modeling memory with historical action trajectories. Drawing on this insight, we represent memory with images. However, directly incorporating historical images into policy input leads to a huge computational burden [48]. To address this problem, we develop an efficient recurrent encoder for modeling visual memory [49], which is explained in Section IV-E. By utilizing historical memory, $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is reduced significantly.
Figure 4: As shown in the sub-figure (a), in the task “pick up a yellow square”, the accurate actions present a multimodal distribution due to the existence of two yellow squares. In the sub-figure (b), if the policy is an unimodal distribution estimator, the action corresponding to the minimum expected error is towards the pink circle, an incorrect moving direction.
III-E Multimodal Distribution Estimation
As described in Section III-B, imitation learning optimizes a policy $\pi$ to capture the probability distribution inherent in demonstration data. Due to the action prediction ambiguity, the expert demonstration distribution is multimodal. Therefore, $\pi$ should be capable of fitting a multimodal distribution.
Recently developed policies mostly regress future actions directly, and the policy network parameters are updated based on loss functions like mean absolute error [8]. Such loss function designs imply these policies regress to unimodal distributions with fixed uncertainties [50]. In other words, these policies are unimodal distribution estimators. In Fig. 4, we show what happens when we employ an unimodal distribution estimator to fit a multimodal distribution. Specifically, the robot needs to pick up a yellow square, and the valid actions present a multimodal distribution due to the existence of two yellow squares. The optimization objective is to minimize the regression error between the multimodal demonstration distribution and unimodal policy prediction, which is formulated in Eq. (4). Similar to the process of mean regression [51], this optimization process leads to an action prediction taking the mean of all valid actions (the mean expected value of this multimodal distribution). As visualized in Fig. 4 (b), the robot moves towards an incorrect target, the middle position between these two yellow squares.
To handle this problem, we propose to model the policy output as Gaussian Mixture Distribution [52] with $K$ components, which can naturally represent multimodal distributions. What the policy predicts is the means, variances, and weights of these Gaussian components, denoted as $\mu_{k}$, $\sigma^{2}_{k}$, and $\alpha_{k}$ for the $k_{\rm th}$ component, respectively. Thus,
$\pi(a_{t}|o_{t},\tau)=\sum\limits_{k=1}^{K}\alpha_{k}\mathcal{N}(\mu_{k},\sigma^{2}_{k}).$ (6)
The component weights $\{\alpha_{k}\}_{k=1}^{K}$ satisfy $\sum\limits_{k=1}^{K}\alpha_{k}=1$ and $\alpha_{k}\geq 0$ for all $k\in\{1,2,\cdots,K\}$. We update the policy following the optimization objective in Eq. (3). In this way, our developed policy is able to fit complex demonstration data with multimodal distributions.
Summarily, the content in this subsection ensures that we can approximate the multimodal demonstration distribution $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$ accurately, and the analysis from Section III-B to Section III-D enables a policy to learn $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$ with fewer demonstrations. Guided by these insights, we develop a generalizable manipulation system GEM, which is elaborated in Section IV.
IV Generalizable Entropy-based Manipulation
In this section, we introduce our developed dynamic object manipulation system GEM. GEM is devised following the theoretical framework in Section III. Only using hundreds of manual demonstrations and thousands of programmatic demonstrations collected from simulated environments, GEM can generalize to diverse unseen target objects, moving dynamics, background environments, and robot embodiments.
Figure 5: Our built hardware platform and remote teleoperation handle. Specifically, Sub-figure (a) illustrates the real robot environment for studying dynamic object manipulation. GEM controls the robot arm to perform diverse tasks given the observations from the external and wrist RGB-D cameras. Sub-figure (b) shows our designed remote handle for teleoperating the robot arm to collect demonstration data.
IV-A Hardware Setup
To study the dynamic object manipulation problem, we build a hardware platform as shown in Fig. 5. In this platform, objects are driven by a conveyor belt at an adjustable velocity. A robot arm with a gripper is mounted on a side of this conveyor belt to manipulate the target object. Two RGB-D cameras are deployed to capture environment observations, one on the robot arm (the wrist camera) and the other on a fixed third-person perspective bracket (the external camera). GEM controls the robot arm to manipulate objects based on the observations from these two cameras.
As illustrated in Fig. 5 (b), a teleoperation handle is designed to teleoperate the robot arm in Fig. 5 (a). Specifically, we control the motion of the robot arm end-effector based on the relative pose change of the handle derived from the SLAM camera. The opening and closing of the gripper is regulated through the trigger pull degree. A battery is mounted on the handle to power all the devices in this handle. The control signals from the SLAM camera and trigger are wirelessly transmitted to the robot arm by the main control board, thereby enabling flexible remote teleoperation. The teleoperation handle is only utilized to collect demonstration data for training the learnable network, which we explain in Section IV-E. The handle is not used when testing the method.
Figure 6: The overall pipeline of GEM. GEM primarily consists of two blocks, the object-centric geometry observation and hybrid action control. In object-centric geometry observation, we exclude the many factors irrelevant to predicting actions from policy observation, and thus the observation entropy is reduced. In hybrid action control, we combine classic visual servo control with learnable network control. In the learnable network control, we design the recurrent probabilistic network, which models historical vision observations as memory. In this way, the action prediction entropy is decreased. Only using hundreds of demonstrations for training, GEM can be generalized to diverse unseen scenes. In this figure, the pink modules are implemented based on learnable neural networks and the yellow modules are non-learnable algorithms.
IV-B GEM Overall Framework
The overall framework of GEM is illustrated in Fig. 6. It can be observed that GEM comprises two blocks, i.e., object-centric geometry perception and hybrid action control. The object-centric geometry perception block extracts 3D point clouds of the target objects in the external and wrist sensor views. The design of this block follows the principles summarized in Section III-C, which points out the importance of excluding factors irrelevant to predicting actions from observations. In this way, the observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is decreased and less data is required to achieve strong generalization.
In hybrid action control, we combine a classic visual servo control algorithm [1] and an imitation learning based neural network control algorithm to compute the robot end-effector trajectories. Specifically, the visual servo control algorithm is responsible for tracking the target object to manipulate, and the neural network control algorithm (the recurrent probabilistic network) generates manipulation actions of different skills according to input skill type instructions. The robot end-effector action is the sum of the actions produced by these two algorithms. This design allows the recurrent probabilistic network not need to consider the moving dynamics (like velocity and acceleration) of the target object, further reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. In addition, the developed recurrent probabilistic network can model long-term historical memory, thereby decreasing the action prediction entropy $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$. The output actions of the recurrent probabilistic network are represented in Gaussian Mixture Distributions as described in Section III-E. Hence, the manipulation multimodal ambiguity problem is addressed.
To enable GEM to adapt to different robot hand embodiments, the hybrid action control block predicts the translation and orientation of the robot end-effector. Based on this prediction, we calculate the corresponding robot joint rotations via inverse kinematics [53], and these rotations are utilized to operate the movements of robot arms. This strategy allows for compatibility with various robot embodiments.
IV-C Object-Centric Geometry Perception
This subsection describes a perception module that produces observations with low $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. To design this module, we first identify the redundant factors in observations. If these factors do not provide necessary information for predicting actions, we can exclude them from observations to reduce $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ while maintaining manipulation performance. Fig. 2 summarizes the primary factors affecting $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. By analyzing them, we first observe that successful manipulation mainly depends on the geometric structures of objects rather than their appearances. However, most embodied networks take images as input [46], which conveys visual appearance directly but lacks explicit 3D geometry. To handle the missing geometry, we employ uncolored 3D points as policy input. These 3D points can be easily accessed via the RGB-D cameras shown in Fig. 5. Besides, some recent deep learning models can produce promising dense depth maps from monocular or binocular images [54, 55], which facilitates obtaining depth. In addition, 3D points provide representation less affected by sensor parameters. This is because 3D points exist in the real world and are not represented in the pixel coordinate system.
Since environment background does not affect manipulation interaction, we adopt a lightweight instance segmentation model [56] to segment points of concerned objects and remove background points. Such an instance segmentation model is readily obtainable due to the availability of many open-world segmentation models [57]. Notably, multiple objects could be observed in a single frame. For example, in the “pick up the spoon out of the cup” task illustrated in Fig. 6, the 3D points of both the spoon and cup are important. In industrial applications, as manipulation procedures are predetermined, we can define which objects to segment in advance. For tasks with unfixed procedures, large language models [26] could be employed to reason about which objects to focus on.
Notably, a manipulation policy needs to classify the points belonging to different objects, like the cup and spoon in the example shown in Fig. 6. Nevertheless, using uncolored 3D points as policy input makes this classification challenging. To address this obstacle, we leverage instance segmentation results to append distinct instance identifiers (IDs) to the points of various objects. With the instance IDs, the object classification problem is addressed. Summarily, combining the above designs, we propose to use uncolored 3D points of concerned objects appended with corresponding instance IDs as the observation representation with low $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
IV-D Hybrid Action Control
This subsection explains how to further reduce $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ by devising the hybrid action control block in Fig. 6. The basic idea is to leverage the complementary strengths of classic control and learnable network control. Classic control algorithms, such as visual servo control [1], excel at stably and rapidly guiding the robot arm to approach and track moving objects. However, they often struggle with collision-rich interactions, such as grasping, inserting, and rotating objects. This is because of the complexity of modeling contact dynamics. On the other hand, learnable network control methods, particularly those grounded in imitation learning, are adept at handling such interactions by capturing the underlying patterns in expert demonstrations. By combining these two approaches, we enable the learnable network to concentrate on generating collision-rich interactions, thereby reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
Figure 7: This figure illustrates how classic control and learnable network control are combined. Specifically, classic control is utilized to approach and track the moving target object. The learnable network control starts generating manipulation actions once stable tracking is achieved. The grippers and target objects marked in gray, dark yellow, and reddish brown correspond to their positions at Moments 1, 2, and 3, respectively. The pink gripper indicates the position that would be achieved using classic control alone at Moment 3.
As described in Section IV-C, the object-centric geometry perception module extracts the 3D points of the target object to manipulate from both the external and wrist sensors. Denote the two groups of 3D points as $\{(x_{i}^{e},y_{i}^{e},z_{i}^{e})\}_{i=1}^{N^{e}}$ and $\{(x_{i}^{w},y_{i}^{w},z_{i}^{w})\}_{i=1}^{N^{w}}$, respectively. The classic control, consisting of an object state estimator and an end-effector tracking planner as shown in Fig. 6, operates the robot hand to track the moving target object based on the external view 3D points. Specifically, the object state estimator first computes the top centroid of the external view 3D points, denoted as $(x_{c}^{e},y_{c}^{e},z_{c}^{e})$, where $x_{c}^{e}=\frac{1}{N_{e}}\sum\limits_{i=1}^{N_{e}}x_{i}^{e}$, $y_{c}^{e}=\frac{1}{N_{e}}\sum\limits_{i=1}^{N_{e}}y_{i}^{e}$, and $z_{c}^{e}=\max\limits_{i=1}^{N_{e}}z_{i}^{e}$ (the positive z-axis points upward). Based on the top centroids at different timestamps, we can estimate the target object velocity $(v_{x},v_{y},v_{z})$ using Gaussian process regression [58] implemented in the object state estimator. If the target object is occluded in the external view camera, its real-time top centroid is also predicted by this estimator based on historical centroid records using Gaussian process regression.
The control target of the end-effector tracking planner is to make the end-effector stably track the motion of the target object. In other words, the relative position offset between the end-effector and the target object should be maintained at the preset distances $(x^{p},y^{p},z^{p})$. In addition, to ensure the wrist RGB-D camera can clearly view the target object, the orientation of the end-effector should be kept at the preset Euler angles $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$. Summarily, the tracking action generated by the end-effector tracking planner comprises three parts, the tracking position $(x_{c}^{e}+x^{p},y_{c}^{e}+y^{p},z_{c}^{e}+z^{p})$, tracking orientation $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$, and tracking velocity $(v_{x},v_{y},v_{z})$.
Based on the tracking action, the end-effector first approaches the target object, as visualized in Moment 1 of Fig. 7. Then, at Moment 2, the end-effector achieves a stable tracking status. In this status, the moving target object appears stationary in the view of the wrist camera. After attaining this status, the learnable network (the recurrent probabilistic network) starts to predict manipulation actions for collision-rich interactions with the target object, as shown in Moment 3. Since the network generates manipulation actions based on the wrist-view 3D points $\{(x_{i}^{w},y_{i}^{w},z_{i}^{w})\}_{i=1}^{N^{w}}$, the original dynamic object manipulation problem is transformed into a static object manipulation problem in the view of the network. Hence, the motion pattern of the target object is excluded from the factors affecting observations, meaning $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is reduced. Besides, this design fixes the relative position between the end-effector and the target object in the beginning of manipulation action generation as $(x^{p},y^{p},z^{p})$, which further decreases $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
The manipulation action predicted by the network includes the position offset $(x^{w},y^{w},z^{w})$ and orientation offset $(\theta^{w}_{x},\theta^{w}_{y},\theta^{w}_{z})$. The end-effector control target output by the hybrid action control block is the sum of the tracking action and manipulation action. Therefore, the end-effector control target consists of the position target $(x_{c}^{e}+x^{p}+x^{w},y_{c}^{e}+y^{p}+y^{w},z_{c}^{e}+z^{p}+z^{w})$, orientation target $(\theta^{p}_{x}+\theta^{w}_{x},\theta^{p}_{y}+\theta^{w}_{y},\theta^{p}_{z}+\theta^{w}_{z})$, and velocity target $(v_{x},v_{y},v_{z})$. Afterwards, the end-effector control target is converted into robot joint rotations based on inverse kinematics. This design enables a single GEM system to be applied to different robot embodiments.
Figure 8: The overall framework of the proposed recurrent probabilistic network, which takes the 3D points of concerned objects as input and includes a memory encoder to manage the historical observation memory.
IV-E Recurrent Probabilistic Network
Section IV-D has described how classic control and learnable network control are combined in the hybrid motion control block. This part further elaborates the structure details of the recurrent probabilistic network used in learnable network control. The overall framework of this network is illustrated in Fig. 8. To concentrate on the geometry of objects rather than appearance, the recurrent probabilistic network takes the wrist view 3D points produced by the object-centric geometry perception block as the main input. For the 3D points perceived at timestamp $t$, it first employs a point feature extractor to encode the 3D points into semantic-rich feature tokens $F_{t}\in\mathbb{R}^{l_{p}\times c}$, where $l_{p}$ denotes the token number and $c$ is the token length. In this work, we model the point feature extractor with Sonata [25], a recent promising point network. However, the input of Sonata should be colored points, while ours are uncolored points appended with instance IDs. To bridge this gap, we feed the instance IDs into Sonata as surrogate color information. Therefore, the points of various object instances in Fig. 8 are visualized with different colors.
The memory encoder shown in Fig. 8 is to fuse $F_{t}$ and the memory feature at timestamp t1t-1, denoted as $M_{t-1}\in\mathbb{R}^{l_{m}\times c}$, into the new memory feature $M_{t}\in\mathbb{R}^{l_{m}\times c}$. To this end, the memory encoder first employs a Transformer decoder [59] to conduct feature interaction between $M_{t-1}$ and $F_{t}$. Specifically, $M_{t-1}$ is used as the query embedding, and $F_{t}$ is the key and value embeddings. The feature output by this Transformer decoder is transformed into the forget score $G_{t}\in\mathbb{R}^{l_{m}\times 1}$ and hidden feature $H_{t}\in\mathbb{R}^{l_{m}\times c}$ by two multilayer perceptron layers (MLPs). Afterwards, in the Mix module visualized in Fig. 8, $M_{t}$ is calculated following:
$M_{t}=\sigma(G_{t})\odot H_{t}+(1-\sigma(G_{t}))\odot M_{t-1},$ (7)
where $\sigma(\cdot)$ represents the Sigmoid function. At the beginning of manipulation, the memory feature is initialized as a zero embedding. Taking these designs, the historical memory is updated in an auto-regressive style, and $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is decreased.
Then, $M_{t}$, $F_{t}$, and a skill type embedding $S_{t}\in\mathbb{R}^{1\times c}$ are concatenated and input to three Transformer decoders. $S_{t}$ is a learnable token specifying which skill to perform, and the supported skills include Pick, Place, Rotate, and Insert. The feature interaction process in decoders is similar to the implementation in ACT [8], where a chunk of $N_{a}$ action tokens is utilized to conduct feature interaction with the input feature. Then, the action tokens are decoded as the action probability distributions of future $N_{a}$ timestamps and a binary finish flag. As described in Section III-E, the action probability distributions are represented as the means, variances, and component weights of Gaussian mixture distribution. The finish flag indicates whether the execution of the manipulation skill has been completed. If completed, the network stops generating actions until the next time of performing a skill. In training, the network is updated following Eq. (3) as described in Section III-E. During testing, the action with the maximum probability is selected for execution.
Figure 9: This figure illustrates our simulated dynamic object manipulation environment. (a) An overview of the environment, where diverse objects are moving on a conveyor belt. (b) We collect 3D assets in 30 representative geometric structures to collect demonstrations. (c) The Pick skill. (d) The Put skill. (e) The Rotate skill. (f) The Insert skill.
IV-F Efficient Demonstration Collection
Combining the aforementioned techniques, we derive a low-entropy manipulation system, which exhibits stronger generalization with less demonstration data. In this part, we introduce how the demonstration data is efficiently collected to support various manipulation skills within a single system.
Demonstration data can be collected in both simulated environments and the real world. Data collection in simulated environments is usually more efficient, but the visual fidelity of existing simulators is still limited. This limitation restricts existing simulated data generation. Thankfully, our network takes uncolored 3D points as input, avoiding the visual fidelity problem. Therefore, our method can generalize to real-world applications even though only using simulated data.
Nevertheless, there is no public simulated environment for dynamic object manipulation. To bridge this gap, we build the first such environment based on Isaac Gym [32]. As shown in Fig. 9 (a), the environment layout is similar to the real platform in Fig. 5. As mentioned in Section IV-C, geometric structure information of objects is essential for successful manipulation. To enable our method to generalize to objects in diverse geometries, we gather 3D assets of 30 representative geometric structures from the GSO dataset [60] to produce demonstration data. These 3D assets are visualized in Fig. 9 (b).
This work supports the four common skills including Pick, Place, Rotate, and Insert. These four skills are shown in Fig. 9 (c)-(f), respectively. Specifically, Pick requires the robot to pick up the moving object from the conveyor belt. Place needs the robot to put an object into a moving container on the conveyor belt. Rotate demands the robot to rotate the moving object to a target orientation. The target orientation in this work is prompted to the network by providing a reference observation, which is the point cloud of the manipulated object in the goal 3D orientation from the perspective of the network. Insert involves inserting an object into a container with a precisely sized opening, ensuring the object is aligned at the specified pose. The Insert task illustrated in Fig. 9 (f) only allows two millimeters of position error.
During gathering demonstrations, all the 30 assets can be used to collect Pick data. For Put, the 12 small-sized assets highlighted by a green box in Fig. 9 (b) are put into the 5 huge-sized containers marked by a yellow box. In Rotate, the 6 assets lacking rotational symmetry are utilized to collect data. For Insert, due to the size matching requirement between the insert object and the container, we only utilize the assets shown in Fig. 9 (f). Notably, the 3D assets are colored with random pure colors instead of being rendered with more realistic appearances. This is because our method takes uncolored point clouds as input, so there is no need to spend computational resources on rendering appearances.
Among these four skills, Pick requires human knowledge to decide reasonable grasp positions. For example, in Fig. 9 (c), the robot should grasp the hammer handle rather than the hammer head. Due to this requirement, the Pick data should be collected through human teleoperation. Specifically, we use our designed teleoperation handle shown in Fig. 5 (b) to collect 20 demonstrations for every 3D asset. In this way, 600 Pick demonstrations are obtained. Different from Pick, the demonstration data of the other three skills can be generated automatically through preprogrammed manipulation procedures. This is because we have the true state information of all objects in simulated environments. Hence, we build an automatic pipeline to generate data for these three skills without human intervention. A total of 2000 demonstrations are derived for each of these three skills.
V Experimental Verifications
V-A Experimental Settings
V-A1 Implementation Details
We train the recurrent probabilistic network using the 6,600 demonstrations collected in Section IV-F for 10K optimization iterations. The training takes about 14 hours on a single RTX4090 GPU. The optimizer is AdamW [61] and the learning rate is 1e41e-4. The training batch size is 32. The image resolutions of the two RGB-D cameras are both $640\times 480$. The action chunk size $N_{a}$, memory token number $l_{m}$, and feature channel number $c$ are set to 20, 200, and 512, respectively. GEM controls the robot at a frequency of 20 HZ.
V-A2 Testing Protocols
We test GEM in both the simulated environment in Fig. 9 and the real platform in Fig. 5. The primary evaluation metric is the task completion success rate. In Pick, the objects in various geometric structures are successively conveyed on the conveyor belt. The success condition for Pick is the gripper successfully picks the moving object up. In Put and Insert, an object and a container are moving on a conveyor belt. Success is defined as the robot arm placing the object inside the container. For Rotate, the robot hand grasps and rotates the object on the conveyor belt to the target orientation. Success is achieved if the yaw, pitch, and roll errors are all within 5 degrees. The default moving speed of the conveyor belt is 0.1 m/s. To mitigate the influence of random errors, we conduct each experiment for 100 times to compute the average success rate.
Figure 10: This figure illustrates the 14 selected objects in diverse geometric structures for out-of-domain manipulation evaluation. All the assets are used for Pick. In Put, the small-sized assets highlighted in a green box are required to be put in the containers marked in yellow. The ten 3D assets lack of rotational symmetry are employed in Rotate.
Figure 11: The 18 objects in diverse geometries that are used to conduct the real-world manipulation experiments. The left 9 objects can be used for collect demonstrations and are used in in-domain evaluation. The right 9 objects are for out-of-domain validation. All the objects can be used in Pick. For Put, one of the small-sized objects highlighted in the green box is put in one of the containers marked in the yellow box. The USB adapter and spoon are employed in Rotate. The objects in the orange box are for Insert, which also only allows two millimeters of manipulation error.
V-B Method Comparison
This subsection compares GEM with previous manipulation methods in dynamic object manipulation. Although the classic kinematics-based methods can also handle dynamic object manipulation after precise system identification and careful parameter tuning, they demand special algorithm designs for different tasks and testing environments. This task-dependent characteristic violates the focus of this work, how to improve the generalization of a manipulation system. Therefore, we mainly compare GEM with the recent learning-based methods, which present better generalization.
The compared methods include AnyGrasp [37], ConvMLP[62], RVT-2 [63], ACT [8], Diffusion Policy (DP) [46], OpenVLA [11], and PI0 [12]. Among them, AnyGrasp can generate grasp pose candidates based on the 3D points of the target object. Then, the robot arm selects the grasp pose with the highest confidence and grasps the target based on kinematic control. Since AnyGrasp only predicts the grasp pose of each object for once, the grasp is executed in an open-loop manner. RVT-2 is a representative two-stage method that takes RGB-D information as input and processes this information based on point cloud reconstruction. The inputs of the other compared methods are the wrist and external view images. Among them, ConvMLP, ACT, and DP are lightweight neural networks based on image observations. OpenVLA and PI0 are two influential vision-language-action (VLA) models. As OpenVLA and PI0 contain billions of parameters, their inference speeds are limited.
TABLE I: Success rate and speed comparison with previous methods in simulation.
[…]
AnyGrasp 0.41 - - - - 0.43 - - - - - 4.35
ConvMLP 0.06 0.19 0.05 0.00 0.08 0.00 0.08 0.00 0.00 0.02 0.05 17.54
ConvMLP* 0.28 0.48 0.19 0.01 0.24 0.06 0.35 0.02 0.00 0.11 0.18 13.16
RVT-2 0.08 0.24 0.03 0.00 0.09 0.00 0.14 0.00 0.00 0.04 0.07 31.48
RVT-2* 0.47 0.62 0.15 0.27 0.38 0.18 0.50 0.00 0.12 0.20 0.29 19.69
ACT 0.14 0.29 0.10 0.00 0.13 0.05 0.22 0.00 0.00 0.07 0.10 43.48
ACT* 0.62 0.75 0.53 0.44 0.59 0.26 0.65 0.08 0.27 0.32 0.46 23.70
DP 0.09 0.24 0.07 0.00 0.10 0.00 0.17 0.00 0.00 0.04 0.07 27.32
DP* 0.55 0.70 0.46 0.36 0.52 0.18 0.54 0.05 0.23 0.25 0.39 17.86
OpenVLA 0.11 0.26 0.08 0.00 0.11 0.07 0.24 0.01 0.00 0.08 0.10 1.61
OpenVLA* 0.64 0.74 0.46 0.35 0.55 0.35 0.68 0.06 0.29 0.35 0.45 1.58
PI0 0.23 0.36 0.18 0.00 0.19 0.12 0.29 0.10 0.00 0.13 0.16 1.97
PI0* 0.68 0.81 0.59 0.65 0.68 0.50 0.75 0.14 0.43 0.46 0.57 1.92
GEM (Ours) 0.88 0.98 0.83 0.80 0.87 0.77 0.91 0.70 0.72 0.78 0.83 23.81
TABLE II: Success rate with previous methods in real world tasks.
[…]
AnyGrasp 0.12 - - - -
ConvMLP 0.00 0.00 0.00 0.00 0.00
ConvMLP* 0.06 0.38 0.02 0.00 0.12
RVT-2 0.02 0.15 0.00 0.00 0.04
RVT-2* 0.17 0.59 0.08 0.04 0.22
ACT 0.08 0.23 0.05 0.00 0.09
ACT* 0.53 0.76 0.40 0.35 0.51
DP 0.05 0.17 0.01 0.00 0.06
DP* 0.46 0.70 0.32 0.25 0.43
OpenVLA 0.00 0.00 0.00 0.00 0.00
OpenVLA* 0.29 0.35 0.14 0.00 0.20
PI0 0.00 0.00 0.00 0.00 0.00
PI0* 0.37 0.58 0.21 0.00 0.29
GEM (Ours) 0.92 0.96 0.79 0.74 0.85
V-B1 Simulated Comparison
This part compares GEM with the aforementioned counterparts in the simulation tasks described in Section V-A. Since AnyGrasp is for generating grasp poses, we only test it in Pick tasks. The algorithms in the hybrid action control module are utilized to help AnyGrasp handle the moving dynamics of objects. The other compared methods are trained using the same demonstrations as our recurrent probabilistic network. Because the compared models have no designs for dynamic object manipulation, we test them in two modes, the sole mode and hybrid mode. The sole mode means the robot movement is solely controlled by a network output. In the hybrid mode, the network output is combined with visual servo tracking like our developed hybrid action control module. In this way, the original dynamic object manipulation problem is converted into a static problem, which favors the performances of the compared methods.
As mentioned in Section IV-F, we collect 30 diverse assets to generate demonstrations and train the compared methods except AnyGrasp. To verify their out-of-domain generalization on unseen geometries, we additionally select 14 new assets, as shown in Fig. 10, to conduct experiments. The out-of-domain Insert validation is conducted using blocks similar to the ones shown in Fig. 9 but we randomly change the block sizes.
The success rates and inference speeds of all the methods are presented in Table I. Specifically, the $2_{\rm nd}-6_{\rm th}$ columns are the success rates of in-domain evaluations based on the assets shown in Fig. 9 (b). The $7_{\rm th}-11_{\rm th}$ columns correspond to the out-of-domain success rates of using the assets in Fig. 10. Avg_in, Avg_out, and Avg refer to the average success rates of in-domain tests, out-of-domain tests, and all tests, respectively. The inference speeds reported in the $13_{\rm th}$ column are the model inference frequencies tested on an RTX4090 GPU. An asterisk (*) following the method name indicates that the method is validated in the hybrid mode. The hybrid mode increases the inference time compared to the sole mode (denoted without the asterisk) because of the additional adoption of an instance segmentation model.
According to the results, we can find that GEM achieves the best success rates in performing all the skills. This is partly because of the efficient structure of the recurrent probabilistic network on perceiving object geometries based on points. In addition, it is observed that the methods in the hybrid mode significantly outperform the ones in the sole mode, which is attributed to that our hybrid action control module alleviates the manipulation challenge. For AnyGrasp, we find it generates promising pick poses. However, it works in an open-loop manner, which restricts its effectiveness. Though RVT-2 uses depth information like our method, its network architecture is inefficient in geometry information exploitation, resulting in its unsatisfactory success rates. Comparing the results of the VLA models and other lightweight networks like ACT, we observe VLAs generally present better out-of-domain generalization thanks to they are pre-trained on numerous data. However, the speeds of VLAs are slower. The slow inference does not affect the success rates significantly in simulated environments because the simulation is paused during model inference, but it harms performances in real tasks significantly.
V-B2 Real Robot Comparison
This study compares GEM with the aforementioned counterparts in real-world tasks using the experiment platform introduced in Fig 5. The objects illustrated in Fig. 11, which are in diverse geometries and sizes, are selected to conduct this comparison. Specifically, we use the left nine objects in Fig. 11 to collect 20 demonstrations for every skill. Then, the total of 80 demonstrations are utilized to fine-tune the policies that have been trained using the simulation data. Finally, employing the same nine objects, we test their in-domain performances in the hybrid mode.
We do not compare the out-of-domain results in this experiment because we find the generalization of the compared methods is generally limited. This is because the color differences between in-domain and out-of-domain objects disturb their action prediction. In the simulation, all the objects are rendered with pure colors, so the compared methods generalize successfully to some extent according to the results in Table I. Nevertheless, in real-world validation, the appearances between in-domain and out-of-domain objects are quite different. Differently, GEM can directly generalize to real world-tasks even though trained utilizing only simulated data, according to the sim-to-real results in Section V-E. This result highlights the advantage of using uncolored 3D points.
The real robot comparison result is presented in Table II, where we can find that GEM presents high success rates after tuning using only 80 demonstrations. Besides, the results of the compared methods in the hybrid mode significantly outperform the results in the sole mode, indicating the value of combining classic control and learning algorithms. Moreover, we can find that although PI0 behaves better than ACT in simulated environments, its success rates are lower in real robot experiments. This is caused by the slow inference of PI0. Therefore, fast inference speed is important for models used in dynamic object manipulation.
Figure 12: Subfigure (a) illustrates the MazeNav task, where a robot is initialized in a random grid of the $5\times 5$ maze and needs to find a path to reach the yellow star. The yellow star is always in the botton-right corner of the maze. Subfigure (b) visualizes the action of the expert policy in every grid.
V-C Theory Verification
This subsection validates the entropy analysis framework described in Section III, i.e., how the changes of observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and action prediction entropy $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ affect the target distribution approximation precision. As the entropy of robotic manipulation is too complex to quantify, we design a toy task named MazeNav to study this problem. Specifically, as shown in Fig. 12 (a), a robot is initialized in a random grid of the $5\times 5$ maze and needs to find a path to reach the destination marked by a yellow star. The destination is always in the bottom-right corner of the maze. The action space includes moving up, down, left, and right. The robot is not allowed to move beyond the maze boundaries.
As illustrated in Fig. 12 (b), a deterministic expert policy is built to generate demonstrations. Based on the demonstrations, a student policy is to estimate the expert policy action distribution by calculating the frequency of each action taken in each state observation. If an action is never performed given a state, the student policy assigns equal possibilities to all candidate actions in this observation state.
Besides the robot position, the state observation of the student policy includes an integer variable $n_{m}$ randomly sampled from $1,N_{m}]$. As a random variable, $n_{m}$ is irrelevant to action prediction. We can change $N_{m}$ to adjust $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. To alter $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, we introduce a threshold variable $\eta\in(0,1)$. During collecting demonstrations, if a variable sampled between $(0,1)$ is smaller than $\eta$, the action is randomly selected from the action space rather than decided by the expert policy. Hence, by increasing $\eta$, $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is enhanced.
Figure 13: This figure illustrates the relation between $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and the target distribution approximation precision. $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is adjusted by changing $N_{m}$. The approximation precision is measured by the KL divergence between the expert distribution and student distribution.
Figure 14: This figure illustrates the relation between $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ and the target distribution approximation precision. $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is adjusted by changing $\eta$. The approximation precision is measured by the KL divergence between the expert distribution and student distribution.
V-C1 Observation Entropy Verification
By adjusting $N_{m}$ and the demonstration trajectory number, this part studies the relation between $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and the expert distribution approximation precision. Specifically, $\eta$ is fixed as 0, meaning that there is no action prediction ambiguity. To study the approximation performance under various $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ values, we repeat this experiment by setting $N_{m}$ to 1, 3, and 5 separately. In each experiment, the expert policy generates 10, 20, 30, 40, and 50 demonstration trajectories to update the student policy, respectively. After the update, the expert distribution approximation precision is measured by the KL divergence between the expert and student action distributions.
To eliminate the influence of random factors, we set different seeds and repeat each experiment five times to calculate the standard deviation. The obtained KL divergence values and standard deviations under different $N_{m}$ and demonstration trajectory numbers are visualized in Fig. 13. We can observe that increasing the demonstration number improves the approximation performance and is helpful in reducing the standard deviation. In addition, by increasing $N_{m}$, which means larger $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, the approximation precision deteriorates when using the same number of demonstrations. These observations are consistent with the theoretical analysis in Section III-C.
V-C2 Action Prediction Entropy Verification
This part explores the relation between $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ and the expert distribution approximation precision. Specifically, $N_{m}$ is fixed as 1, indicating that the policy observation does not include redundant information. We set $\eta$ to 0.0, 0.3, and 0.6 separately to adjust $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$. Similar to the previous experiment, we update the policy using 10, 20, 30, 40, and 50 demonstration trajectories, respectively. The experiment of each hyper-parameter is repeated five times to compute the mean KL divergence and standard deviation. The experimental results of different $\eta$ and demonstration numbers are shown in Fig. 14.
According to the results, we can find that a larger $\eta$, which indicates bigger $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, leads to poorer distribution estimation precision. In addition, by incorporating more demonstrations, the student policy can still estimate a promising expert policy distribution even though $\eta$ is high. These results confirm the theoretical analysis in Section III-D.
Figure 15: This experiment suggests that the student policy can perfectly estimate the action choices of the expert policy if sufficient demonstrations are provided, even though the demonstrations are very noisy (high $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$) and contain much irrelevant observation (high $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$). Specifically, the subfigure (a) visualizes the expert policy. Subfigures (b)-(d) show the student policies updated with 50, 200, and 1000 demonstrations, respectively. The student policy using 1000 demonstrations perfectly replicates the expert policy.
V-C3 Effect of Data Scaling Up
This experiment aims to validate the hypothesis in Section III-B that the optimal actions can be precisely approximated by scaling up expert demonstrations. To study this hypothesis, we set $N_{m}$ to 10, which leads to high $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$. Besides, $\eta$ is 0.9, meaning the robot performs random actions in 90% of the steps. This results in a large $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$.
We update the student policy with 50, 200, and 1000 demonstrations. Then, we visualize the student policy actions corresponding to the biggest probabilities under various states in Fig. 15 (b)-(d). We can find that even though the demonstration data is highly noisy, the student policy still perfectly replicates the expert policy using 1000 demonstrations. This result confirms the hypothesis that the optimal actions can be precisely approximated with sufficient demonstrations.
TABLE III: Comparison between point and imgae observations.
[…]
Image $\times$ 0.75 0.60 0.64 0.33
Image $\checkmark$ 0.41 0.28 0.38 0.14
Point $\times$ 0.79 0.71 0.73 0.66
Point $\checkmark$ 0.87 0.78 0.85 0.74
Figure 16: This figure visualizes some instance segmentation cases during manipulation, where we can find that the segmentation masks are precise.
V-D Component Analysis
This subsection ablates the various designs in GEM. Due to the space limit, when simultaneously reporting the in-domain and out-of-domain results in simulated and real environments, we report the average success rates of all skills.
V-D1 Analysis on Point based Representation
Different from previous works commonly taking 2D images as input, our method utilizes 3D points. This design is based on the intuition that accurate manipulation relies on the geometry of objects, and 3D points better deliver such information. This experiment is to validate this intuition. The results are presented in Table III, where the first two rows of results correspond to the models based on images, and the last two rows are the models using points. When the input is in the image modality, the feature extractor of the recurrent probabilistic network is changed to DINOv2 [64], a widely adopted network due to its discriminative representation. The second column indicates whether the background regions are removed based on instance segmentation results. “Sim_in”, “Sim_out”, “Real_in”, and “Real_out” refer to the average success rates in the settings of in-domain simulation, out-of-domain simulation, in-domain real world, out-of-domain real world, respectively.
According to the results, we can first find that the models taking 3D points as input outperform the ones with images, which confirms our intuition. Secondly, it is observed that the operation of removing background regions favors point based models but harms the performance of image based models. This is because when the input is an image, the model demands background texture to infer the depth of the target object. By contrast, the model does not need background since the object position has been known from 3D points.
V-D2 Analysis on Instance Segmentation
The results in Table III have revealed how the operation of removing background based on instance segmentation affects manipulation success rates. This part analyzes the efficiency and robustness of our employed instance segmentation model, YOLOv11. Since the official YOLOv11 only supports segmenting objects in the categories defined in the COCO dataset [65], we need to collect a few additional instance segmentation data to enable YOLOv11 to handle our concerned categories. Thankfully, the development of SAM2 [57] makes this data collection process efficient. Specifically, based on SAM2, we only need to record a video for every concerned category of objects and click a point on each object to generate segmentation masks. In this way, obtaining tuning data totally takes several minutes. The inference speed of YOLOv11 evaluated on an RTX4090 GPU is more than 100FPS, meeting the real-time requirement.
Afterwards, to evaluate the segmentation precision of the tuned YOLOv11, we collect a dataset consisting of 1800 images for the objects shown in Fig. 11 (200 frames per object). Then, we utilize this dataset to validate the segmentation precision of YOLOv11. Following the testing protocol in COCO, YOLOv11 achieves a Mask AP of 94.5% in this dataset, suggesting that the segmentation precision YOLOv11 satisfies the manipulation task demand. Furthermore, we visualize some challenging cases in Fig. 16 during manipulation. It is found that the segmentation quality is promising.
Figure 17: This figure visualizes the tracking dynamics of the end-effector under different target velocities. We can find that the end-effector stably tracks objects moving at 0.2 m/s but fails to keep up with objects moving at 0.6 m/s.
TABLE IV: Manipulation performance in various moving speeds.
[…]
0.05 0.92 0.96 0.79 0.74
0.10 0.90 0.97 0.80 0.71
0.25 0.81 0.92 0.73 0.48
0.50 0.00 0.00 0.00 0.00
V-D3 Analysis on Visual Servo tracking
In GEM, we utilize the visual servo tracking module to convert the dynamic object manipulation problem into static object manipulation in the view of the recurrent probabilistic network. This study analyzes the tracking stability and velocity upper limit in real-world manipulation. Specifically, we adjust the conveyor belt to various speeds and make the robot hand track a moving object based on the visual servo tracking module. It is observed that the maximum stable tracking speed in simulation is approximately 0.25 m/s, and it reaches about 0.3 m/s in the real world. In Fig. 17, we visualize the dynamic curves of the control target signal and corresponding end-effector position. We can find that the end-effector approaches the moving object fast and tracks it stably under the velocity of 0.2 m/s. However, the end-effector fails to track the object when the velocity is 0.6 m/s due to its speed limit.
Afterwards, we validate the real-world success rates of GEM in the settings of different moving speeds. The results are reported in Table IV. Comparing the $1_{\rm st}$ and $2_{\rm nd}$ rows of results, we can observe that the success rates are similar when the conveyor belt runs at a low speed. When the speed is 0.25 m/s, the success rates drop due to stricter manipulation precision requirements. When the speed arrives at 0.50 m/s, the robot hand cannot perform any manipulation task because it cannot keep pace with the movement of the target object.
TABLE V: Analysis on the effect of the memory encoder.
[…]
$\times$ 0.72 0.60 0.69 0.55
$\checkmark$ 0.87 0.78 0.85 0.74
V-D4 Analysis on Memory Modeling
As discussed in Section III-D, enabling a network to access historical memory is important for reducing action prediction ambiguity. To this end, we devise the memory encoder shown in Fig. 8. In experiments, we find that without this encoder, the robot arm could oscillate at the initial position of manipulation without approaching the target, leading to task failure. This phenomenon is more frequent when a single network learns multiple skills. This is because the policy fails to distinguish whether the manipulation operation has not started or has been completed. Our designed memory encoder can effectively address this problem by enabling the network to know which actions have been performed. We compare the performances of GEM with and without this memory encoder and report the results in Table V. According to the results, we can observe that the memory encoder significantly boosts success rates.
To more thoroughly analyze how long a sequence the memory encoder can memorize, we design a toy experiment of reciting a random digit sequence. Specifically, we update the structure of this memory encoder to support accepting a single digit as input and predicting the next digit per step. Then, we randomly generate a sequence of $n$ digits and train the memory encoder to recite this sequence. After training, we test how many digits can be correctly recited. After repeating this experiment by setting different values of $n$, we find that the memory encoder can correctly recite a very long sequence like 1000 digits after sufficient epochs of training. Such a memory capability is strong enough for existing embodied policies, because current manipulation tasks usually involve at most tens of times of policy inference.
Figure 18: This figure visualizes the simulated environment for analyzing manipulation multimodal ambiguity. During collecting demonstrations in this environment, we teleoperate the robot hand to randomly pick up one of these two blocks and put it in the green container.
TABLE VI: Analysis on network output.
[…]
Direct 0.81 0.73 0.79 0.67
Probabilistic 0.87 0.78 0.85 0.74
V-D5 Analysis on Network Output
As explained in Section III-E, the output of the recurrent probabilistic network is a Gaussian mixture distribution to represent the predicted action probabilistic distribution. This is different from previous works that directly regress the actions to execute [18, 46]. This part analyzes the benefits of our probabilistic output.
We first compare the performances of networks based on direct action prediction and probabilistic action prediction. The results are reported in Table VI. We can observe that the recurrent probabilistic network based on probabilistic action prediction outperforms the one using direct action prediction, because the multimodal ambiguity problem is alleviated.
Notably, we argue that the results in Table VI do not sufficiently reveal the value of probabilistic action prediction. This is because the recurrent probabilistic network only observes a single target object due to the object-centric geometry perception block, and the actions in the collected demonstrations are always to manipulate this single object. There is no case like the one shown in Fig. 4, where multiple target objects appear simultaneously in the network observation.
To bridge this gap, we develop a simulated environment as shown in Fig. 18. We use the teleoperation handle shown in Fig. 5 (b) to control the robot hand to collect 50 demonstrations, where we randomly pick up one of these two blocks and put it into the green container. To study multi-target ambiguity, we remove the object-centric geometry perception block in GEM and make the recurrent probabilistic network take images as input. The feature extractor is changed to DINOv2. After training using the collected demonstrations, we compare the success rates of the networks based on direct action prediction and probabilistic action prediction. We find that the one based on direct action prediction makes the robot end-effector hover at the midpoint between the two blocks, leading to a success rate of nearly 0%. By contrast, the network based on probabilistic action prediction achieves a success rate of 87%. These results demonstrate the importance of adopting our proposed probabilistic action prediction.
TABLE VII: Analysis on hybrid action control.
[…]
$\times$ 0.35 0.28 0.20 0.16
$\checkmark$ 0.87 0.78 0.85 0.74
V-D6 Analysis on Hybrid Action Control
The hybrid action control block combines visual servo control and learnable network control to conduct dynamic object manipulation. This experiment studies how this strategy contributes to success rates compared with the common design, i.e., solely based on learnable network control. The performance comparison results between GEMs with and without using the visual servo control are presented in Table VII. We can find that the visual servo control, which converts the original dynamic object manipulation problem into static manipulation, is critical for the overall task completion success rates.
TABLE VIII: Analysis on the effect of simulated demonstrations.
[…]
$\times$ 0.28 0.45 0.15 0.08
$\checkmark$ 0.92 0.96 0.79 0.74
V-D7 Analysis on Data Efficiency
As explained in Section IV-F, since the recurrent probabilistic network takes uncolored 3D points as input, the observations in simulators become similar to real-world observations. Therefore, the simulated data can directly benefit real-world performances. Correspondingly, we collect 6,600 demonstrations in our built simulation environment. The policy trained using these simulated demonstrations should significantly favor the real-world results. To verify this issue, we compare the real-world performances of the GEM policies with and without being pre-trained with these simulated demonstrations. Both the two policies are tuned using the 80 real-world demonstrations described in Section V-B before the real world evaluation.
The experimental results are presented in Table VIII. We can find that the simulated data boosts the success rates of various skills by large margins. According to the first row of results, only utilizing the 80 real-world demonstrations is insufficient to derive a promising policy, and collecting more real-world demonstrations is labor-intensive. Thankfully, simulated demonstrations can tackle this data scarcity problem because simulated data generation can be efficient.
V-E Generalization Analysis
The various designs in GEM are to improve the generalization of this manipulation system with minimal cost. This subsection conducts experiments to validate the generalization capabilities of GEM from different perspectives.
Figure 19: To study the generalization performance of GEM across different environment backgrounds, we randomly augment the appearance of the conveyor belt, and these generated appearances are unseen in demonstrations. Two augmented samples are visualized in this figure.
TABLE IX: Environment generalization study in simulation.
[…]
Seen 0.88 0.98 0.83 0.80
Unseen 0.91 0.96 0.83 0.81
V-E1 Generalization across Environment Backgrounds
Thanks to the rapid progress of instance segmentation [57, 66], obtaining a segmentation model meeting manipulation precision requirements has become highly cost-effective. Therefore, we design the object-centric geometry perception block in GEM that produces uncolored 3D points appended with instance IDs as the input to the recurrent probabilistic network. In this way, regardless of changes in the environment background, the input to the network remains similar. Therefore, our method should present strong generalization across different environmental backgrounds.
To validate this characteristic, we test the manipulation performance of GEM in different unseen backgrounds by changing the appearance of the conveyor belt. Since changing conveyor belts is expensive to realize in the real world, we conduct this experiment in simulation. As shown in Fig. 19, we randomly augment the colors of the conveyor belt. Since all demonstrations are collected using the conveyor belt illustrated in Fig. 9 (a), these augmented conveyor belts are unseen in demonstrations. The manipulation success rates of GEM in the original seen environment and the augmented unseen environments are presented in Table. IX. We can observe that the performances of GEM are similar in seen and unseen environments, which is consistent with the expectation. In addition, the subsequent sim-to-real experiment in Table XIII indicates that the GEM trained using only simulated data can generalize to real-world manipulation despite the appearance gap between the simulated and real-world environments.
Figure 20: This figure illustrates how we evaluate the motion pattern generalization capability of GEM in simulated and real environments. Different from the linear motion in demonstrations, in subfigure (a), the object moves following an S-shaped trajectory. In subfigure (b), the object is manually held and moves along a random curved trajectory.
TABLE X: Generalization study on motion patterns.
[…]
Simulated Linear 0.88 0.98 0.83 0.80
Simulated S-curve 0.91 0.97 0.80 0.77
Real Linear 0.92 0.96 0.79 0.74
Real Random curve 0.88 0.98 0.75 0.66
V-E2 Generalization across Motion Patterns
As the visual servo control in the hybrid action control block transforms the original dynamic object manipulation problem into a static manipulation problem, the GEM policy should can generalize to manipulating objects in diverse motion patterns, e.g., various moving velocities and trajectories. The experimental results in Table IV have confirmed that the manipulation success rates remain similar when the conveyor belt speed is adjusted but does not surpass the robot hand tracking velocity limit. In this experiment, we further study this problem by altering the object moving trajectories. Specifically, in simulation, we program the objects to move on the conveyor belt following an S-shaped trajectory as illustrated in Fig. 20 (a). In real-world evaluation, we manually hold a container and put the object to manipulate in this container, moving it along a random curved trajectory, as shown in Fig. 20 (b).
We report the experiment results in Table X. Comparing the performance between linear and curved object movement trajectories, we can conclude that making objects move along an unknown curved trajectory does not harm the manipulation success rates significantly. Therefore, GEM is robust to different motion patterns of moving objects.
Figure 21: This figure visualizes the four kinds of robot hands to conduct cross-embodiment evaluation in the simulated and real environments.
TABLE XI: Generalization study on object geometries.
[…]
Seen 0.92 0.96 0.79 0.74 0.85
Unseen 0.80 0.90 0.67 0.59 0.74
V-E3 Generalization across Object Geometries
By collecting demonstrations based on objects in diverse geometries and training the recurrent probabilistic network using these demonstrations, GEM can generalize to different geometries. The experiment results in Table I suggest that GEM can perform various manipulation skills on unseen objects with promising success rates. Table II reports the performance of GEM on manipulating seen objects in real-world tasks, where the results are similar to the ones in simulated environments. In this experiment, we further reveal the real-world success rates on manipulating the right nine unseen objects in Fig. 11. The recurrent probabilistic network validated in this experiment is the same as the one in Table II, which is first pre-trained using the 6,600 simulated demonstrations and then fine-tuned with 20 real-world demonstrations per skill.
The experiment results are reported in Table XI. We can find that although the network is only tuned with 20 demonstrations per skill, it still generalizes well to unseen objects in diverse geometries. In Section V-F, we show that if we fine-tune the network using 100 Pick demonstrations, it achieves a success rate of over 97% in an unseen application environment.
V-E4 Generalization across Robot Embodiments
Thanks to the visual servo control, the end-effector moves to a fixed track position relative to the target object before that the recurrent probabilistic network begins generating manipulation actions. Additionally, the robot hand is controlled by the predicted end-effector relative offset based on inverse kinematics. Therefore, GEM is not restricted within a single robot embodiment. To evaluate the robot embodiment generalization, we conduct experiments using four kinds of robot hands in simulated and real-world environments, as shown in Fig. 21. In the simulation, the policy trained using the simulated demonstrations collected from Franka Panda is directly applied to the UR10e robot hand. In the real-world experiment, the policy tuned based on the demonstrations collected with Piper is deployed on the ARX5 robot hand to conduct manipulation.
The experiment results are reported in Table XII. It is observed that employing a robot hand to collect demonstrations and adopting another robot hand to conduct manipulation evaluation achieves similar performance compared with the setting where the demonstration collection and validation are based on the same robot hand. Hence, the robot embodiment generalization capability of GEM is confirmed.
Figure 22: Subfigure (a) shows the real application, where the robot hand needs to collect spoons from the trays moving on a conveyor belt. Subfigures (b)-(d) illustrate some challenging cases, i.e., spoons with food residue, multiple spoons to pick up within a short time window, and spoons stacked together.
TABLE XII: Generalization study on robot embodiments.
[…]
Franka Franka 0.88 0.98 0.83 0.80
Franka UR10e 0.91 0.97 0.85 0.79
Piper Piper 0.92 0.96 0.79 0.74
Piper ARX5 0.90 0.95 0.77 0.69
TABLE XIII: Analysis on sim-to-real generalization.
[…]
$\times$ 0.76 0.83 0.55 0.48
$\checkmark$ 0.92 0.96 0.79 0.74
V-E5 Sim-to-Real Generalization
The input to the recurrent probabilistic network is uncolored 3D points, which do not exhibit significant differences between simulated and real environments. Therefore, GEM solely based on simulated demonstrations can generalize to the real world to perform operations. To validate this characteristic, we report the real-world manipulation success rates of GEMs with and without tuning in Table XIII. According to the results, we can observe that the GEM trained using only simulated data can manipulate objects in the real world, though its success rates are lower than the GEM tuned with 20 demonstrations per skill. Through analysis, we find this is caused by two reasons. First of all, without tuning, the test objects are all unseen in demonstrations, which increases the manipulation challenge. Secondly, in a robot control system, there is a control execution delay between the input signal and the system response. This control execution delays are different between simulated and real robots, further enhancing the difficulty.
V-F Real Application Deployment
Different from lab environments, it is often infeasible to collect demonstrations in real application scenes because of the uninterrupted normal workflows and strict confidentiality requirements. Therefore, if a policy is to be deployed, it should be capable of operating in unseen environments with consistently high success rates, handling unexpected object poses and lighting changes. The aforementioned techniques proposed in this work help GEM meet these demands.
To evaluate the robustness of GEM, we deploy it in a canteen kitchen to conduct tableware collection. Specifically, after dining, canteen customers place their used tableware on a conveyor belt, which transports these items to the kitchen. The manipulation system needs to identify and pick out the spoons from the tableware and place them into a separate container, as shown in Fig. 22 (a). The spoons are placed in varying poses and may shift during transport. In addition, as depicted in Fig. 22 (b)-(d), other challenges include food residue obscuring spoon pose recognition, the robot arm needing to grasp and place multiple spoons within tight time windows, and spoons stacked together, etc.
Since onsite data collection is infeasible in this canteen, we teleoperate 100 demonstrations using the manipulation platform shown in Fig. 5. Then, we utilize these 100 demonstrations to fine-tune the GEM policy that has been pre-trained with the 6,600 simulated data. Finally, we deploy the fine-tuned GEM in this canteen to conduct spoon collection. We evaluate GEM over a seven-day period, processing more than ten thousand spoons, and measure a final manipulation success rate of 97.2%. This marks the first real application of imitation learning in dynamic object manipulation.
VI Conclusion
This paper has first developed an entropy-based theoretical framework to analyze the data requirements of imitation learning. Based on this framework, we have proposed GEM that reduces observation complexity and action prediction ambiguity to achieve strong generalization. Through extensive experiments, GEM has successfully generalized across diverse environment backgrounds, motion patterns, object geometries, and robot embodiments, achieving a success rate of over 97% in a real-world application without using any in-scene demonstration. These results have demonstrated that our theoretical insights effectively guide the design of an efficient manipulation system, offering a practical solution for industrial automation with reduced data costs.

### v2

@@
-Train Once, Deploy Anywhere: Realize Data-Efficient Dynamic Object Manipulation
+Sim-to-Real Dynamic Object Manipulation on Conveyor Systems via Optimization Path Shaping
 Abstract
-Realizing generalizable dynamic object manipulation is important for enhancing manufacturing efficiency, as it eliminates specialized engineering for various scenarios.
+Realizing generalizable dynamic object manipulation on conveyor systems is important for enhancing manufacturing efficiency, as it eliminates specialized engineering for different scenarios.
 To this end, imitation learning emerges as a promising paradigm, leveraging expert demonstrations to teach a policy manipulation skills.
 Although the generalization of an imitation learning policy can be improved by increasing demonstrations, demonstration collection is labor-intensive.
-To address this problem, this paper investigates whether strong generalization in dynamic object manipulation is achievable with only a few demonstrations.
-Specifically, we develop an entropy-based theoretical framework to quantify the optimization of imitation learning.
-Based on this framework, we propose a system named Generalizable Entropy-based Manipulation (GEM).
-Extensive experiments in simulated and real tasks demonstrate that GEM can generalize across diverse environment backgrounds, robot embodiments, motion dynamics, and object geometries.
-Notably, GEM has been deployed in a real canteen for tableware collection.
-Without any in-scene demonstration, it achieves a success rate of over 97% across more than 10,000 operations.
-Index Terms:
-Dynamic object manipulation, imitation learning, information entropy, manufacturing efficiency.
+Besides, public dynamic object manipulation data is scarce.
+In this work, we address this data scarcity problem via generating demonstrations in a simulator.
+A significant challenge of using simulated data lies in the appearance gap between simulated and real-world observations.
+To tackle this challenge, we propose Geometry-Enhanced Model (GEM), which employs our designed appearance noise annealing strategy to shape the policy optimization path, thereby prioritizing the geometry information in observations.
+Extensive experiments in simulated and real-world tasks demonstrate that GEM can generalize across environment backgrounds, robot embodiments, motion dynamics, and object geometries.
+Notably, GEM is deployed in a real canteen for tableware collection.
+Without test-scene data, GEM achieves a success rate of over 97% across more than 10,000 operations.
 I Introduction
 Robotic manipulation of dynamic objects has broad applications in industrial manufacturing, particularly in conveyor-based assembly lines where robots handle transported products [38, 37].
-Compared with the extensively studied static object manipulation problem [3, 4], dynamic object manipulation poses more challenges due to its demand for more time-efficient motion planning and stricter operational precision.
-Addressing these challenges is important for enhancing the efficiency and flexibility of industrial automation systems, thereby enabling more economical manufacturing.
+Compared with the extensively studied static object manipulation problem [23, 18], dynamic object manipulation on conveyor systems poses more challenges due to its demand for more time-efficient motion planning and stricter operational precision.
+Addressing these challenges is important for enhancing the efficiency and flexibility of industrial automation systems, thereby enabling more economical manufacturing and sustainable production [43].
 Currently, the methods employed in real-world production lines are primarily based on classic motion planning approaches [46].
@@
 Additionally, even in static object manipulation, the generalization of the policies trained using these large-scale datasets is still limited [24, 4], implying the volume of demonstration is still far from sufficient.
-Therefore, although realizing generalizable dynamic object manipulation is of huge importance, collecting demonstrations is costly and there are few public data resources [13].
-Given these facts, this work studies whether strong generalization in dynamic object manipulation can be achieved without relying on numerous demonstrations.
-To address this question, we delve into the optimization essence of imitation learning, framing it as a distribution approximation problem.
-Building upon this formulation, we propose an information entropy based theoretical framework to quantify the amount of data required to achieve generalizable manipulation.
-From this theoretical framework, we identify two key aspects: first, reducing the complexity of redundant information in observations, and second, decreasing the ambiguity in action prediction.
+Therefore, although realizing generalizable dynamic object manipulation on conveyor systems is important, collecting demonstrations is costly, and public data resource is scarce [58].
+Given these facts, this work studies whether we can generate demonstrations in a simulator and achieve sim-to-real dynamic object manipulation.
+Since all real environment states are available in a simulator, we can generate numerous demonstrations via motion planning efficiently.
+Built upon this idea, we build a dynamic object manipulation simulator based on Isaac Gym [31], as visualized in Fig.
+1.
+Nevertheless, although the simulator handles the collision dynamics between objects well, the appearance gap between simulated and real-world environments is significant.
 Figure 1: Primarily using demonstrations collected from a simulator, our method can generalize across diverse environment backgrounds, robot embodiments, motion dynamics, and object geometries.
 Our method has been reliably deployed in a real canteen to conduct tableware collection.
-Without relying on any demonstration collected in this canteen, our method achieves a success rate of over 97% in seven consecutive days of operation, performing more than 10,000 tableware collection operations.
-According to these two key aspects, we develop a manipulation system named Generalizable Entropy-based Manipulation (GEM).
-GEM consists of two blocks, i.e., object-centric geometry perception and hybrid action control.
-In object-centric geometry perception, we devise a representation that emphasizes the geometric structures of objects while minimizing the impact of visual appearance, thereby reducing redundant details irrelevant to manipulation.
-Within hybrid action control, imitation learning is combined with visual servo control to manage the moving dynamics of objects.
-Additionally, an efficient memory modeling strategy is proposed in the imitation learning algorithm to alleviate action prediction ambiguity.
-In experiments, we implement a single GEM to support four common skills in dynamic object manipulation, i.e., Pick, Put, Rotate, and Insert.
-Extensive experiments are conducted in both simulators and the real world to validate the effectiveness of GEM.
-The experiments suggest that GEM presents promising generalization across diverse environment backgrounds, robot embodiments, motion dynamics, object geometries, and in sim-to-real settings.
+Without using demonstrations collected in this canteen, our method achieves a success rate of over 97% in seven consecutive days of operation, performing more than 10,000 tableware collection operations.
+To bridge this appearance gap, our core insight is that the information in policy observation can be decomposed into geometric structure and visual appearance.
+Importantly, manipulation primarily relies on geometric structure, which is consistent between simulated and real-world environments.
+Based on this insight, we propose a policy named Geometry-Enhanced Model (GEM), which takes colored 3D points from RGB-D cameras as input.
+In GEM, we propose an appearance noise annealing strategy to shape the policy optimization path, reducing the reliance of the policy on appearance while enhancing its focus on geometry.
+In this way, sim-to-real generalization is enabled.
+Moreover, we propose a control strategy that disentangles manipulation into tracking actions and interaction actions.
+This strategy empowers GEM to manipulate objects moving in unseen velocities.
+We conduct extensive experiments in both simulated and real-world tasks to validate the effectiveness of GEM.
+The experiments suggest that GEM presents promising generalization across diverse environment backgrounds, robot embodiments, motion dynamics, and object geometries.
 Especially, as illustrated in Fig.
-1, GEM has been successfully deployed in a real canteen to perform tableware collection.
+1, GEM has been successfully deployed in a real canteen to collect tableware.
 Without using demonstrations collected in this canteen, GEM achieves a 97.2% success rate over seven days of testing and more than 10,000 operations.
-Summarily, the contributions of this work are as follows:
-1) We develop an entropy-based framework to analyze the optimization of imitation learning, guiding the design of an efficient manipulation system with limited demonstrations.
-2) We propose GEM, a system combining object-centric geometry perception and hybrid action control to achieve strong generalization in dynamic object manipulation.
-3) We demonstrate the first real-world application of imitation learning in dynamic object manipulation, achieving over 97% success rate without in-scene demonstrations.
+Summarily, the technical contributions of this work are summarized as follows:
+1) We build the first simulation environment for generating dynamic object manipulation data on conveyor systems.
+2) We propose a neural policy GEM, which realizes sim-to-real generalization by prioritizing the exploitation of geometry information while reducing the reliance on appearance.
+3) We design a control strategy that disentangles manipulation into tracking and interaction actions, enabling GEM to manipulate dynamic objects of diverse motion patterns.
 II Related Works
@@
 However, manipulation often involves contact-rich interactions with objects, and accurately modeling these interactions is difficult due to the highly variable nature of contact forces [19].
-Moreover, every manipulation task often requires a unique set of models and parameters, making these traditional methods difficult to generalize across different objects, environments, and robot embodiments.
+Moreover, every single manipulation task often requires a unique set of models and parameters, making these traditional methods difficult to generalize across different objects, environments, and robot embodiments.
 This necessitates substantial engineering efforts to adapt the system to new tasks or scenarios [59].
@@
 Specifically, imitation learning trains a policy to predict expert actions based on the observations provided in demonstrations, enabling a robot to mimic expert behavior in similar scenarios.
-The policy is often implemented as a deep learning network [16], which can capture the complex patterns in demonstrations [17].
+The policy is often implemented as a deep learning network [26], which can capture the complex patterns of demonstrations [6].
 The two most commonly employed observation modalities in robotic manipulation are images and 3D points [46].
@@
 Images are easy to obtain, and image-based policies can readily leverage large-scale public data for training [33].
-However, robotic manipulation mainly concerns the geometric structures of objects, while images primarily capture the appearance of scenes, making it challenging to predict appropriate manipulation actions based solely on visual appearance.
-In contrast, point clouds directly represent the geometric structures of scenes, and thus they are more suitable for regressing manipulation actions [21].
+However, robotic manipulation mainly concerns the geometric structures of objects, while images primarily capture the appearance of scenes, making it challenging to predict appropriate manipulation actions solely based on visual appearance.
+In contrast, point clouds directly represent the geometric structures of scenes, and thus they are more suitable for estimating manipulation actions [52].
 Unfortunately, point cloud data is relatively scarce compared with images, resulting in a lack of promising point cloud feature extractors [39].
@@
 Leveraging these datasets, numerous studies design algorithms for training end-to-end manipulation policies, with the expectation that these policies exhibit robust generalization across diverse scenarios [24, 29].
-Furthermore, some works incorporate additional types of data like object detection datasets to further enrich the knowledge base of manipulation policies [12, 28].
+Furthermore, some works incorporate additional types of data like object detection datasets to further enrich the knowledge of manipulation policies [4, 3].
 Nevertheless, though much resource has been invested, the generalization capabilities of the obtained policies in unseen environments are still limited.
@@
 Collecting high-quality demonstration data in the real world is labor-intensive.
-To address this issue, many researchers have explored generating demonstrations in simulated environments [30].
-These efforts aim to develop policies trained using only simulated demonstrations yet can generalize to real-world applications.
-A key first step towards this goal is the development of realistic simulators.
+To address this obstacle, many researchers have explored generating demonstrations in simulated environments [32].
+These efforts aim to develop policies that solely employ simulated demonstrations to train yet can generalize to real-world applications.
+A key step towards this goal is the development of realistic simulators.
 Through substantial engineering efforts, several promising simulators have been built, such as MuJuCo [48], Isaac Gym [31], and Isaac Sim [34].
 Leveraging these simulators, prior works have used reinforcement learning [20] or predefined kinematic procedures [9] to generate demonstrations.
-However, reinforcement learning often struggles to discover successful manipulation trajectories due to its inefficient exploration [36].
-Grasping objects via predefined kinematic procedures typically requires customized workflows for different scenarios, which is time-consuming.
-Additionally, current simulators have limited environmental diversity, restricting the variety of generated data.
-Another line of works focuses on modeling the environment dynamics of manipulation tasks using classic mechanics and then employs optimization techniques to calculate actions that satisfy task constraints [14].
+However, reinforcement learning often struggles to discover appropriate manipulation trajectories due to its inefficient exploration [47].
+Grasping objects via kinematic procedures typically requires customized workflows for different scenarios, which is time-consuming.
+Additionally, current simulators have limited environmental diversity, restricting the variety of the generated data.
+Another line of literature focuses on modeling the environment dynamics of manipulation tasks using classic mechanics and then employs optimization techniques to calculate actions that satisfy task constraints [19].
 The observation and action pairs can then be utilized to train deep neural networks [12, 49].
 This strategy effectively addresses the slow speed of optimization by distilling the action search process into a more efficient neural network [16].
-Nevertheless, for complex manipulation tasks, the computation burden is too heavy to search for satisfactory actions, making it difficult to generate effective demonstrations.
-To tackle this problem, some works collect demonstrations in simulators through teleoperating simulated robots [40].
+Nevertheless, for complex manipulation tasks, the computation burden for searching promising action trajectories could be heavy, making it difficult to generate effective demonstrations.
+To tackle this problem, some works collect demonstrations in simulators through teleoperating robots [61].
 Although this method still requires significant human effort, it reduces hardware setup costs compared to real-world data collection.
-III Entropy Reduction Theoretical Framework
-This section develops a theoretical framework by analyzing the information entropy of imitation learning, providing logical support for developing a generalizable manipulation system with only a few demonstrations.
-III-A Robotic Manipulation Problem Formulation
+III Problem Analysis
+This section outlines the policy design principle of this work, providing a theoretical analysis of how we shape the optimization path of an imitation learning policy.
+Specifically, Section III-A formulates the policy optimization objective of robotic manipulation.
+Next, Section III-B discusses policy observation information and explains why 3D geometry is more important than 2D appearance in embodied tasks.
+Subsequently, Section III-C introduces the core process of imitation learning.
+Finally, Section III-D elaborates the concept of optimization path shaping in imitation learning, which is the theoretical foundation of this work.
+III-A Problem Formulation
 Robotic manipulation involves controlling a robot arm to interact with objects in the environment to achieve a specific task.
@@
 In a specific task target $\tau\in\mathcal{T}$ of robotic manipulation, there usually exist numerous accurate action trajectories that can lead the robot from the initial state $s_{0}$ to the desired goal state $\iota$.
-We denote all the accurate trajectories as a set $\mathcal{C}$ and each trajectory is represented as a sequence of states and actions $\{s_{0},a_{0},\cdots,s_{T-1},a_{T-1},s_{T}\}$, where $s_{T}=\iota$.
+We denote all the accurate trajectories as a set $\mathcal{C}$ and each trajectory can be represented as a sequence of states and actions $\{s_{0},a_{0},\cdots,s_{T-1},a_{T-1},s_{T}\}$, where $s_{T}=\iota$.
 To enable the robot to automatically complete the target task $\tau$ without human intervention, researchers devise algorithms to find an optimal policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the probability of reaching the target state $\iota$ following the transition probability distribution $\mathcal{P}$.
@@
 This formulation explicitly connects policy optimization with implicit distribution estimation of $\mathcal{C}$.
-III-B Entropy Analysis of Imitation Learning
 However, in robotic manipulation, a policy $\pi$ usually does not have access to the true state $s_{t}$ at timestamp $t$.
@@
 The policy $\pi$ is often modeled as a deep neural network [26].
-Many recent manipulation works derive $\pi$ based on imitation learning [8], which is typically formulated as a supervised learning problem.
-Imitation learning requires access to a set of expert demonstrations as training data.
+Figure 2: The 3D geometry of objects usually plays a more important role than 2D appearance in scheduling manipulation actions.
+As shown, although the left and middle objects share similar color and texture, the appropriate gripper poses to pick them up are different.
+By contrast, though the middle and right objects belong to different object categories with distinct appearances, they can be picked up with a similar gripper pose.
+III-B Policy Observation Analysis
+According to the aforementioned formulation, the environment observation $o_{t}$ provides the information that is utilized by $\pi$ to schedule actions.
+The two most commonly employed observation modalities for recent manipulation policies are images captured by cameras [13] and point clouds returned by depth sensors [56].
+Images directly describe environment appearance, allowing $\pi$ to more accurately identify object categories and materials.
+Point clouds focus more on object geometry, enabling a policy to determine manipulation poses more precisely.
+Both the appearance and geometry information are valuable for successful manipulation.
+To combine their benefits, this work adopts colored point clouds obtained from RGB-D cameras as policy input, enabling the policy to perceive both appearance and geometry information appropriately.
+Notably, in the tasks of dynamic object manipulation on conveyor belts, geometry is more critical than appearance.
+Without color information, a policy can still identify the moving objects on a conveyor belt in most cases.
+Moreover, the policy can often determine accurate manipulation actions based only on geometry.
+For instance, as illustrated in Fig.
+2, while the left and middle objects share similar color and texture, their accurate gripper poses for grasping are different.
+In contrast, despite belonging to distinct object categories with unique appearances, the middle and right objects can be picked up using similar gripper poses.
+Therefore, a policy should focus more on exploiting the geometry information.
+In addition, as visualized in Fig.
+1, while existing simulators can effectively replicate the geometric structure of objects, the rendered appearance still exhibits a significant gap compared to the real world.
+Thankfully, as indicated by the above analysis, successful manipulation in our explored tasks primarily requires accurate geometry information.
+Hence, in this work, we expect $\pi$ to concentrate on 3D geometry rather than 2D appearance.
+To achieve this, we design strategies to shape the optimization dynamics in imitation learning.
+For a clear explanation of these strategies, we first introduce the core process of imitation learning in Section III-C.
+III-C Imitation Learning
+Imitation learning is typically formulated as a supervised learning problem [60].
+It requires access to a set of expert demonstrations as training data.
 Assume there are $N_{d}$ demonstration sequences, the expert demonstration set is denoted as $\mathcal{C}_{d}=\{o_{0:T}^{i},\tilde{a}_{0:T}^{i}\}_{i=1}^{N_{d}}$, where $o_{t}^{i}$ and $\tilde{a}_{t}^{i}$ are the observation and expert action at timestamp $t$ of the $i_{\rm th}$ demonstration.
-The core idea of imitation learning is employing $\mathcal{C}_{d}$ as a surrogate of $\mathcal{C}$.
+The core idea of imitation learning is employing $\mathcal{C}_{d}$ as a surrogate of the successful manipulation set $\mathcal{C}$.
 Thus, the objective of imitation learning can be expressed as maximizing the likelihood of expert demonstrations under the policy-induced distribution:
@@
 Inspired by the recent breakthroughs in research fields like large language models [1] and image generation [42], it is believed that the distribution of $\mathcal{C}$ can be precisely approximated by scaling up the volume of expert demonstration data $\mathcal{C}_{d}$ [24, 35].
-Nevertheless, collecting diverse demonstrations is expensive in robotic manipulation [42].
-Although huge efforts have been invested in building large-scale manipulation datasets [9, 10], the available quantity of demonstrations is far from sufficient compared with fields like large language models [43].
-Currently, the policies trained on these datasets only serve as pre-trained network weights, still requiring task-specific fine-tuning for each downstream application [27, 42].
-Thus, exploring strategies to reduce the demonstration data requirement of imitation learning is important.
-To conduct this exploration, we first specify that the optimization objective of imitation learning is to approximate the joint observation-action distribution $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$, which describes how to achieve successful manipulation.
-For notational convenience, we represent the sets of observations and actions in $\mathcal{C}$ as $\mathcal{O}_{\mathcal{C}}$ and $\mathcal{A}_{\mathcal{C}}$.
-To explore how to reduce the demonstration demand while maintaining policy performance, we need a theoretical framework to quantify the data requirement for approximating $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$.
-In information theory, the entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}},\mathcal{A}_{\mathcal{C}})$ serves as a measure of the amount of information contained within a distribution, reflecting the inherent uncertainty or complexity of data [44].
-A higher entropy indicates the distribution contains more information and requires more demonstrations to accurately approximate.
-Furthermore, we decompose $\mathcal{H}(\mathcal{O}_{\mathcal{C}},\mathcal{A}_{\mathcal{C}})$ via the chain rule:
-$\mathcal{H}(\mathcal{O}_{\mathcal{C}},\mathcal{A}_{\mathcal{C}})=\mathcal{H}(\mathcal{O}_{\mathcal{C}})+\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}}).$ (5)
-This decomposition reveals that the information entropy of imitation learning consists of two parts, the observation complexity $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and the action prediction ambiguity $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$.
-By reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, the demonstration volume requirement can be alleviated.
-In addition, we expect the manipulation performance not to compromise after reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$.
-To figure out how to realize this goal, it should be first noted that although there are numerous manipulation trajectories meeting constraints, $\pi$ only needs to predict one of them to complete the task.
-Therefore, when reducing the entropy, the probability of selecting a suitable action leading to successful manipulation should not be decreased.
-To ensure this, (i) When lessening the observation complexity $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, we only remove redundant observation information irrelevant to inferring accurate actions.
-(ii) During reducing the action prediction ambiguity $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, the policy should increase the likelihood of predicting a successful trajectory.
-Figure 2: This figure illustrates the primary factors affecting the observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-The factors include environment, robot, sensor, and task instruction.
-These factors can be further divided.
-III-C Observation Complexity Reduction
-To alleviate the demonstration data demand, we first study how to reduce the observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, which measures the complexity of the observation set in $\mathcal{C}$.
-Specifically, we expect to reduce the redundant information in observations that is irrelevant to inferring actions, because removing such information does not hurt action prediction precision.
-Notably, which kinds of information are irrelevant to actions are not deterministic.
-They are influenced by how the manipulation system is designed.
-For instance, if a policy $\pi$ directly predicts the robot hand end-effector movement and then utilizes this movement to control joint rotations based on inverse kinematics, the joint angle observation is unnecessary for inferring actions when the end-effector position observation has been provided.
-Therefore, the key to reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is devising a manipulation system making most factors irrelevant to inferring actions.
-Then, these irrelevant factors can be excluded from observations, thereby reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-To more thoroughly analyze this problem, we need to know $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is determined by which factors.
-We illustrate the primary factors in Fig.
-2.
-These main factors contributing to $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ can be categorized into four classes, i.e., environment information, robot proprioception, sensor configuration, and task instruction.
-Among them, the environment information includes the target object to manipulate, scene background, and ambient conditions (such as lighting brightness and air humidity).
-These pieces of information can be further divided, like the target object information can be classified into its physical geometry and visual appearance.
-Differently, robot proprioception mainly describes the type of robot employed (like Franka Panda or UR5e), where the robot is located in the environment, and the action space (like end-effector control or joint control).
-Sensor configuration specifies what kinds of sensors are adopted (like RGB camera and depth sensor), how the sensors are deployed in the environment (the number of sensors and their poses), and the parameters of these sensors.
-A task instruction specifies the task target.
-It can be represented in different modalities, including text [11], image [40], voice [45], etc.
-Summarily, the above four classes of factors play distinct roles in shaping the complexity of the observation set $\mathcal{O}_{\mathcal{C}}$.
-To reduce $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, we need to design a system making most factors irrelevant to predicting actions.
-We explain how such a specific system is devised in Section IV.
-Figure 3: The sub-figure (a)-(c) illustrate the procedures that a robot hand presses a button.
-We can find that if the policy input does not include historical memory and moving velocity, the policy observations are the same between sub-figure (a) and (c), in which the robot hand is approaching the button and moving away from the button, respectively.
-III-D Action Prediction Ambiguity Reduction
-In this subsection, we discuss how to reduce $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ in Eq.
-(5), which represents the action prediction ambiguity given observations.
-This ambiguity is mainly caused by the multimodal nature of robotic manipulation, meaning there exist multiple valid action trajectories given the same observation [46].
-For example, in the task of grasping a block, a robot arm can complete the task by approaching the block from either the left or the right side.
-Moreover, the ambiguity becomes more pronounced when the observation information is insufficient.
-For instance, Fig.
-3 illustrates an example that a robot hand is pressing a button.
-In Fig.
-3 (a)-(c), the robot hand is approaching the button, pressing the button, and departing from the button, respectively.
-If the policy input only includes the current visual observation and robot joint angles, the observations are the same between Fig.
-3 (a) and (c), while the accurate end-effector moving directions are different.
-Hence, larger action prediction ambiguity not only increases demonstration demand, but also raises the risk of making incorrect decisions.
-To handle this problem, more comprehensive information supporting accurate action prediction should be provided.
-According to the example in Fig.
-3, incorporating historical memory into policy observation is effective, especially for complex and long-chained manipulation tasks.
-This is because historical memory enables the policy to know which manipulation steps have been completed.
-However, previous works often find incorporating historical memory degrades success rates [8].
-One key reason is that these works choose to model memory using historical action trajectories.
-As analyzed in VIP [40], when the attribute to predict (the future action trajectory in this case) and a part of the policy input (the historical action trajectory) fall in the same modality, the policy often chooses to overlook the input information in other modalities (like images).
-This is because it is easier to predict future actions based on historical actions than information in different modalities.
-A learnable network tends to choose the easiest way of regressing future actions while overlooking the exploitation of information in other modalities.
-This problem is termed as shortcut learning and has been well analyzed in previous literature [47].
-An effective strategy for handling the shortcut learning problem is preventing using input in the same modality as output.
-In this case, it means avoiding modeling memory with historical action trajectories.
-Drawing on this insight, we represent memory with images.
-However, directly incorporating historical images into policy input leads to a huge computational burden [48].
-To address this problem, we develop an efficient recurrent encoder for modeling visual memory [49], which is explained in Section IV-E.
-By utilizing historical memory, $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is reduced significantly.
-Figure 4: As shown in the sub-figure (a), in the task “pick up a yellow square”, the accurate actions present a multimodal distribution due to the existence of two yellow squares.
-In the sub-figure (b), if the policy is an unimodal distribution estimator, the action corresponding to the minimum expected error is towards the pink circle, an incorrect moving direction.
-III-E Multimodal Distribution Estimation
-As described in Section III-B, imitation learning optimizes a policy $\pi$ to capture the probability distribution inherent in demonstration data.
-Due to the action prediction ambiguity, the expert demonstration distribution is multimodal.
-Therefore, $\pi$ should be capable of fitting a multimodal distribution.
-Recently developed policies mostly regress future actions directly, and the policy network parameters are updated based on loss functions like mean absolute error [8].
-Such loss function designs imply these policies regress to unimodal distributions with fixed uncertainties [50].
-In other words, these policies are unimodal distribution estimators.
-In Fig.
-4, we show what happens when we employ an unimodal distribution estimator to fit a multimodal distribution.
-Specifically, the robot needs to pick up a yellow square, and the valid actions present a multimodal distribution due to the existence of two yellow squares.
-The optimization objective is to minimize the regression error between the multimodal demonstration distribution and unimodal policy prediction, which is formulated in Eq.
-(4).
-Similar to the process of mean regression [51], this optimization process leads to an action prediction taking the mean of all valid actions (the mean expected value of this multimodal distribution).
-As visualized in Fig.
-4 (b), the robot moves towards an incorrect target, the middle position between these two yellow squares.
-To handle this problem, we propose to model the policy output as Gaussian Mixture Distribution [52] with $K$ components, which can naturally represent multimodal distributions.
-What the policy predicts is the means, variances, and weights of these Gaussian components, denoted as $\mu_{k}$, $\sigma^{2}_{k}$, and $\alpha_{k}$ for the $k_{\rm th}$ component, respectively.
-Thus,
-$\pi(a_{t}|o_{t},\tau)=\sum\limits_{k=1}^{K}\alpha_{k}\mathcal{N}(\mu_{k},\sigma^{2}_{k}).$ (6)
-The component weights $\{\alpha_{k}\}_{k=1}^{K}$ satisfy $\sum\limits_{k=1}^{K}\alpha_{k}=1$ and $\alpha_{k}\geq 0$ for all $k\in\{1,2,\cdots,K\}$.
-We update the policy following the optimization objective in Eq.
-(3).
-In this way, our developed policy is able to fit complex demonstration data with multimodal distributions.
-Summarily, the content in this subsection ensures that we can approximate the multimodal demonstration distribution $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$ accurately, and the analysis from Section III-B to Section III-D enables a policy to learn $p_{\mathcal{C}}(o_{0:T},a_{0:T-1})$ with fewer demonstrations.
-Guided by these insights, we develop a generalizable manipulation system GEM, which is elaborated in Section IV.
-IV Generalizable Entropy-based Manipulation
-In this section, we introduce our developed dynamic object manipulation system GEM.
-GEM is devised following the theoretical framework in Section III.
-Only using hundreds of manual demonstrations and thousands of programmatic demonstrations collected from simulated environments, GEM can generalize to diverse unseen target objects, moving dynamics, background environments, and robot embodiments.
-Figure 5: Our built hardware platform and remote teleoperation handle.
-Specifically, Sub-figure (a) illustrates the real robot environment for studying dynamic object manipulation.
+III-D Optimization Path Shaping
+The optimization objective of imitation learning is solely to minimize the cost value calculated by the difference measurement function defined in Eq.
+(4) and does not specify which observation feature to exploit.
+As discussed in Section III-B, the observation feature in colored point clouds can be categorized into the appearance and geometry information.
+If an RGB-D policy learns to only use appearance information, its learning process becomes similar to an image-only policy, like ACT [60].
+In this situation, the policy needs to infer the 3D geometry of objects from visual appearance implicitly to estimate accurate manipulation poses [54].
+By contrast, if a policy primarily relies on geometry, it can predict manipulation poses in a more direct and precise manner.
+As discussed in Section III-B, the appearance gap between simulated and real environments motivates this work to shape the policy optimization path to prioritize the geometric information, which is more consistent between simulated and real-world observations.
+To analyze how to achieve this target, we draw on the deep learning theory of feature predictivity established in prior literature [17].
+In this theory, predictivity basically means how useful a feature is for accurately inferring training labels, and a policy tends to pick features with higher predictivity to minimize training loss.
+Since the geometry of an object directly reflects its 3D structure, geometry is inherently of higher predictivity for estimating manipulation actions.
+However, in simulation, scene layout is often not diverse, e.g., objects with specific textures always appear in fixed areas.
+This makes appearance features (e.g., color, texture) accidentally gain unexpected high predictivity.
+Correspondingly, the policy overfits to these scene-specific appearance patterns, using them to predict actions instead of the generalizable geometry.
+To prioritize geometry information and alleviate appearance overfitting, we propose to reduce the predictivity of appearance features by introducing visual noise, thereby shaping the policy optimization path towards geometry.
+However, persistent appearance noise would prevent the policy from leveraging useful appearance cues.
+For instance, in block insertion tasks where a block should be placed into a container, appearance helps distinguish these two objects, and relying solely on geometry could lead to misidentification.
+To resolve this appearance utilization trade-off, we propose appearance noise annealing, a strategy that gradually removes appearance noise during training.
+Specifically, strong appearance noise is applied at the beginning of training to guide the policy to focus on geometry, followed by steady noise reduction until the noise is fully removed.
+Since the policy already learns to use geometry for loss minimization in early training stages, it will keep prioritizing geometry after noise is removed.
+For scenarios where geometry alone is insufficient (e.g., distinguishing closely placed distinct objects), the policy can learn to exploit useful appearance texture after the appearance feature becomes obvious.
+IV Policy Design
+This section presents the proposed manipulation policy, GEM.
+Specifically, Section IV-A introduces the overall framework of GEM.
+Then, Section IV-B and Section IV-C explain the developed appearance noise annealing and manipulation action decomposition strategies separately.
+Afterwards, Section IV-D describes the network implementation details in GEM.
+Finally, Section IV-E elaborates how we build simulation environments of dynamic object manipulation on conveyor systems to generate demonstration data efficiently.
+Figure 3: The overall framework of GEM.
+The policy observation is the colored 3D points captured by RGB-D cameras.
+The appearance noise annealing strategy is utilized to perturb the color feature in policy observation.
+Based on this observation, the GEM network predicts a multi-modal action distribution represented as a Gaussian Mixture Model (GMM), a point segmentation mask of the target object to manipulate, and a status flag marking the completion of manipulating one object.
+We generate interaction actions and tracking actions from the network outputs and combine them to control the robot end-effector.
+IV-A Policy Framework
+We illustrate the overall framework of the GEM policy in Fig.
+3.
+To ensure sufficient information in the policy environment observation, we mount two RGB-D cameras in the dynamic object manipulation system.
+Specifically, an external-view camera is installed opposite the robotic arm, with its perspective directed at the conveyor belt.
+Another wrist-view camera is affixed to the wrist of the robotic arm.
+The colored 3D points perceived by these two cameras are merged as a single point cloud to provide environment observation.
+As mentioned in Section III-D, we design an appearance noise annealing strategy that shapes the policy optimization path by applying strong color augmentation to the point cloud at the beginning of training and gradually removing the noise until the training end.
+This strategy guides the policy to focus more on the 3D geometry of objects.
+Next, as shown in Fig.
+3, the augmented point cloud is input to the GEM network, which consists of Transformer encoders, decoders, and a memory module.
+The memory module represents the historical observation information as an implicit embedding and updates it in a recurrent manner.
+Incorporating memory information into the network input helps the network be aware of the completion progress of manipulation.
+The prediction produced by the GEM network includes an action probabilistic distribution, a point segmentation mask, and a status flag as visualized in Fig.
+3.
+The action probabilistic distribution is modeled using a Gaussian Mixture Model (GMM) [41].
+Unlike the popular implementations of recent policies that directly estimate deterministic actions [13], the GEM network predicts a probabilistic distribution over actions, as such an output naturally captures the inherent indeterminism of manipulation decisions.
+The experimental results suggest that this design can effectively alleviate the multi-modal ambiguity in imitation learning and is more efficient than the commonly employed strategies like diffusion process [7] and flow matching [28].
+During inference, we decode the interaction actions from the GMM distribution by selecting the action with the maximum probability value at each timestamp.
+The point segmentation prediction is to segment out the 3D points belonging to the target object.
+This prediction has two benefits, i.e., guiding the policy to focus on the target object and generating tracking actions.
+For the first benefit, we expect the GEM network to concentrate on the 3D geometry of the target object rather than the whole environment, as the environment background has little effect on successful manipulation.
+We can understand this guiding process from the view of optimization path shaping, since it increases the importance of target object points on minimizing training loss.
+The second benefit is we can approximate the 3D centers of target objects based on the segmentation results and make the robot end-effector track the target objects moving in various velocities.
+We realize this dynamic object tracking by decoding the point cloud segmentation results into tracking actions.
+Both the interaction actions and tracking actions represent the relative pose offsets of the robot end-effector.
+By combining them following our developed algorithm based on the predicted status flag, we obtain the manipulation actions utilized to control the rotations of robot joints based on inverse kinematics [25].
+This design of predicting end-effector relative offsets enables the policy to generalize across different robot embodiments.
+Refer to Section IV-C for the details of how we generate and utilize the predicted actions.
+IV-B Appearance Noise Annealing
+Figure 4: In the appearance noise annealing strategy, strong color perturbation (noise ratio is 1.0) is applied to the input point cloud at the initial phase of training.
+As the training progresses, the perturbation is gradually removed.
+As discussed in Section III-D, we develop a simple yet effective strategy to shape the policy optimization path, enabling it to focus on object geometry and generalize to real-world tasks via simulated demonstrations.
+In this strategy, we apply strong color perturbation to the input point cloud at the initial stage of training.
+According to the aforementioned theory of feature predictivity, this operation reduces the feature predictivity of 2D appearance, thereby compelling the GEM network to prioritize 3D geometry in minimizing training loss.
+Then, with the progress of training, the color perturbation is gradually weakened.
+Since the network has already learned to predict manipulation actions using the observed 3D geometric structure of the environment, it does not revert to over-relying on visual appearance after the color perturbation is removed.
+Furthermore, after the visual features become prominent, the network begins to learn how to leverage visual appearance to handle tasks that cannot be adequately addressed by 3D geometry alone, such as distinguishing two adjacent objects.
+In the following, we elaborate on the implementation details of how the color perturbation is applied.
+Denote the input point cloud at timestamp $t$ with $N$ points as $Q^{t}=\{q_{i}^{t}\}_{i=1}^{N}$, where $q_{i}^{t}=(x_{i}^{t},y_{i}^{t},z_{i}^{t},r_{i}^{t},g_{i}^{t},b_{i}^{t})$.
+In $q_{i}^{t}$, $(x_{i}^{t},y_{i}^{t},z_{i}^{t})$ is the point coordinate in the wrist-view camera coordinate system, and $(r_{i}^{t},g_{i}^{t},b_{i}^{t})$ represents the visual color of this point.
+The color channels have been normalized to the interval $[0,1]$, i.e., $(r_{i}^{t},g_{i}^{t},b_{i}^{t})\in[0,1]^{3}$.
+To disturb the color of the point $q_{i}^{t}$, we sample three independent random scalars $\eta_{i,r}^{t},\eta_{i,g}^{t},\eta_{i,b}^{t}\sim\rm{Uniform}(0,1)$ and assemble them into an RGB noise vector denoted as $\eta_{i}^{t}=(\eta_{i,r}^{t},\eta_{i,g}^{t},\eta_{i,b}^{t})\in[0,1]^{3}$.
+To control the strength of the noise perturbation $\eta_{i}^{t}$ applied to $(r_{i}^{t},g_{i}^{t},b_{i}^{t})$, we define a noise ratio factor $\gamma_{\xi}$ that varies with training progress, where $\xi\in[0,1]$ denotes the normalized training progress.
+To ensure a strong perturbation at the beginning of training and make the perturbation gradually vanish as learning proceeds, the variable $\gamma_{\xi}$ is defined as:
+$\gamma_{\xi}=\begin{cases}1,&\xi\in[0.0,0.2),\\ 1-\sigma(k(\xi-m)),&\xi\in[0.2,0.8],\\ 0,&\xi\in(0.8,1.0],\end{cases}$ (5)
+where $\sigma(\cdot)$ is fomulated as $\sigma(x)=1/(1+e^{-x})$, and $k$ and $m$ are set to 10 and 0.5, respectively.
+The curve of $\gamma_{\xi}$ evolving with the training progress is illustrated in Fig.
+4.
+The point color $c_{i}^{t}=(r_{i}^{t},g_{i}^{t},b_{i}^{t})$ is augmented into the perturbed vector $\tilde{c}_{i}^{t}$ following:
+$\tilde{c}_{i}^{t}=\gamma_{\xi}\eta_{i}^{t}+(1-\gamma_{\xi})c_{i}^{t}.$ (6)
+We can observe from Fig.
+4 that the appearance observation of the environment gradually becomes obvious with the progress of training, which satisfies the algorithm design requirement of optimization path shaping.
+The point cloud with the perturbed color attributes $\{\tilde{c}_{i}^{t}\}_{i=1}^{N}$ serves as the direct input to the GEM network.
+If $\gamma_{\xi}$ is 1, the visual feature observed from the point cloud is pure random noise.
+When $\gamma_{\xi}$ becomes 0, the appearance observation is noise-free.
+IV-C Manipulation Acition Decomposition
+As illustrated in Fig.
+3, GEM first decodes the network output into the tracking and interaction actions and then combines them as the manipulation actions.
+Thanks to this strategy, GEM can generalize to manipulating objects moving in diverse unseen velocities, even though the GEM network is trained solely on static object demonstrations.
+This capability makes GEM practical for real-world industrial manipulation on conveyor belts, where the belt speed varies.
+To explain this control strategy clearly, we first describe how to derive the tracking actions in the following.
+Figure 5: This figure illustrates how the predicted tracking actions and interaction actions are integrated to perform dynamic object manipulation.
+Specifically, tracking actions are responsible for approaching and tracking the moving target object.
+The interaction actions begin to take effect once stable tracking is achieved, and they are for conducting collision-rich contact with the target object, such as picking and rotating.
+Based on the point segmentation mask from the GEM network, we extract the spatial coordinates of the 3D points belonging to the target object to manipulate, denoted as $\{(x_{i}^{\tau},y_{i}^{\tau},z_{i}^{\tau})\}_{i=1}^{N^{\tau}}$, where $N^{\tau}$ is the total number of 3D points.
+The top centroid of these points is defined as $(x^{c},y^{c},z^{c})$, where $x^{c}=\frac{1}{N^{\tau}}\sum\limits_{i=1}^{N^{\tau}}x_{i}^{\tau}$, $y_{c}=\frac{1}{N^{\tau}}\sum\limits_{i=1}^{N^{\tau}}y_{i}^{\tau}$, and $z_{c}=\max\limits_{i=1}^{N^{\tau}}z_{i}^{\tau}$ (the positive z-axis points upward).
+By tracking this top centroid across time, we can derive the translational velocity $(v^{\tau}_{x},v^{\tau}_{y},v^{\tau}_{z})$ of the target object with simple estimation algorithms like Gaussian process regression [50].
+The target of tracking actions is to make the end-effector stably track the motion of the target object.
+In other words, the relative position offset between the end-effector and the target object should be maintained at the preset distances $(x^{p},y^{p},z^{p})$.
+In addition, to ensure the wrist RGB-D camera can clearly view the target object, the orientation of the end-effector should be kept at the preset Euler angles $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$.
+Therefore, to maintain the synchronous motion of the robot end-effector and the moving target object in a relatively stationary state, the tracking action should include the tracking position $(x^{c}+x^{p},y^{c}+y^{p},z^{c}+z^{p})$, tracking orientation $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$, and tracking velocity $(v_{x},v_{y},v_{z})$.
+By solely executing the tracking action as the manipulation action, the end-effector first approaches the target object, as visualized at Timestamp 1 of Fig.
+5.
+Then, at Timestamp 2, the end-effector achieves a stable tracking status.
+In this status, the moving target object appears stationary in the view of the wrist camera.
+Upon attaining this stable status, as illustrated at Timestamp 3 of Fig.
+5, the interaction action predicted by the GEM network begins to be sequentially added to the tracking action, which then serves as the composite manipulation action.
+The interaction action includes the position offset $(x^{n},y^{n},z^{n})$ and pose offset $(\theta^{n}_{x},\theta^{n}_{y},\theta^{n}_{z})$.
+In this control strategy, the tracking action alone ensures that the target object is in a relatively stationary state with respect to the robot end-effector.
+This means when the GEM network predicts interaction actions, it behaves as though it were manipulating a stationary object.
+In this way, the influence of various conveyor speeds on generating interaction actions is explicitly decoupled and removed.
+Therefore, this action decomposition strategy enables the GEM network trained using static object manipulation demonstrations to operate objects moving at diverse speeds on conveyor systems.
+Summarily, before the end-effector achieves the stable tracking status, the manipulation action is identical to the tracking action computed from the point segmentation result, and it comprises the position goal $(x^{c}+x^{p},y^{c}+y^{p},z^{c}+z^{p})$, orientation goal $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$, and velocity goal $(v_{x},v_{y},v_{z})$.
+After stable tracking is attained, the manipulation action is the sum of the tracking action and interaction action, including the position goal $(x^{n}+x^{c}+x^{p},y^{n}+y^{c}+y^{p},z^{n}+z^{c}+z^{p})$, orientation goal $(\theta^{n}_{x}+\theta^{p}_{x},\theta^{n}_{y}+\theta^{p}_{y},\theta^{n}_{z}+\theta^{p}_{z})$, and velocity goal $(v_{x},v_{y},v_{z})$.
+The manipulation action is transformed into the control signal of robot joints and end-effector based on inverse kinematics and PID [2].
+Notably, the status flag predicted by the policy indicates whether the manipulation of an object has been completed.
+Once completed, the interaction action is removed from the manipulation action, and the tracking action switches to the next target object.
+IV-D Network Implementation
+Figure 6: The overall structure of the GEM network, which consists of four point encoders, three Transformer decoders, and a memory module.
+The memory module updates the memory representation in a recurrent manner.
+The network outputs include the GMM-based action distribution, point segmentation mask, and a status flag that marks the completion of manipulating one object.
+The overall framework of the GEM network is illustrated in Fig.
+6.
+During training, the network input is the 3D points perturbed by the appearance noise annealing strategy.
+In inference, the input is the points without noise.
+For the 3D points perceived at timestamp $t$, the network encodes the points into a semantic-rich feature embedding $F_{t}\in\mathbb{R}^{l_{p}\times c}$ using four point encoders [52], where $l_{p}$ denotes the token number and $c$ is the token length.
+We initialize the parameters of these encoders with Sonata [51].
+Each token in $F_{t}$ corresponds to a voxel in the 3D physical space.
+In the point segmentation branch, a multilayer perceptron (MLP) layer is applied to transform $F_{t}$ into the segmentation mask by performing binary classification.
+The segmentation branch optimization objective is defined as a binary cross-entropy loss [26].
+Subsequently, the memory module shown in Fig.
+6 fuses $F_{t}$ and the memory feature at timestamp t1t-1, denoted as $M_{t-1}\in\mathbb{R}^{l_{m}\times c}$, into the new memory feature $M_{t}\in\mathbb{R}^{l_{m}\times c}$.
+During this process, the memory encoder first employs a Transformer decoder [5] to conduct feature interaction between $M_{t-1}$ and $F_{t}$.
+Specifically, $M_{t-1}$ serves as the query embedding, and $F_{t}$ is the key and value embeddings.
+The feature output by this Transformer decoder is transformed into the forget score $G_{t}\in\mathbb{R}^{l_{m}\times 1}$ and hidden feature $H_{t}\in\mathbb{R}^{l_{m}\times c}$ by two MLP layers.
+Afterwards, $M_{t}$ is calculated following:
+$M_{t}=\sigma(G_{t})\odot H_{t}+(1-\sigma(G_{t}))\odot M_{t-1},$ (7)
+where $\sigma(\cdot)$ represents the Sigmoid function.
+At the beginning of manipulation, the memory feature is initialized as a zero embedding.
+Incorporating historical memory into the network observation is beneficial for enabling GEM to accurately assess the current manipulation progress, thereby reducing the ambiguity in network optimization.
+Then, $F_{t}$ and $M_{t}$ are concatenated and input to three Transformer decoders to conduct feature interaction with $N_{a}$ action tokens.
+The interaction process in decoders is similar to the implementation of ACT [60], where the $N_{a}$ action tokens correspond to the $N_{a}$ actions scheduled for the subsequent $N_{a}$ timestamps.
+After the interaction in these decoders, we transform these action tokens into the probabilistic distributions of actions.
+This operation is different from recent works that commonly decode action tokens into deterministic action values [24].
+We choose to predict action distributions because this design better captures the multi-modal nature of embodied actions, thereby alleviating the optimization conflict in training caused by decision-making ambiguity.
+For each action token, we decode it into an action distribution represented using GMM with $K$ components.
+Specifically, an MLP layer predicts the means, variances, and weights of these $K$ Gaussian components based on each action token.
+For the $t_{\rm th}$ ($t\in[1,N_{a}]$) action token, denote the mean, variance, and weight of the $k_{\rm th}$ component as $\mu_{t,k}$, $\sigma_{t,k}^{2}$, and $\alpha_{t,k}$, respectively.
+The distribution of the $t_{\rm th}$ scheduled action $p(a_{t})$ can be formulated as:
+$p(a_{t})=\sum\limits_{k=1}^{K}\alpha_{t,k}\mathcal{N}(\mu_{t,k},\sigma^{2}_{t,k}).$ (8)
+The component weights $\{\alpha_{t,k}\}_{k=1}^{K}$ satisfy $\sum\limits_{k=1}^{K}\alpha_{t,k}=1$ and $\alpha_{t,k}\geq 0$ for all $k\in\{1,2,\cdots,K\}$.
+In training, assuming the corresponding expert actions of the $N_{a}$ timestamps are $\{\tilde{a}_{t}\}_{t=1}^{N_{a}}$.
+According to Eq.
+(3), the supervision loss for action prediction is defined as:
+$L_{a}=-\frac{1}{N_{a}}\sum\limits_{t=1}^{N_{a}}\log(\sum\limits_{k=1}^{K}\frac{\alpha_{t,k}}{\sqrt{2\pi\sigma_{t,k}^{2}}}\exp(-\frac{(\tilde{a}_{t}-\mu_{t,k})^{2}}{2\sigma_{t,k}^{2}})).$ (9)
+During inference, the highest-probability action at each of the $N_{a}$ future timestamps is selected, forming an action chunk that is executed in a single rollout.
+Moreover, the GEM network predicts a Boolean status flag that indicates whether the manipulation of the current object is completed.
+As described in Section IV-C, this flag controls when the policy should switch to the next target object.
+In training, the status flag prediction is supervised with a binary cross-entropy loss.
+IV-E Data Generation in Simulation
+[…]
+Combining the aforementioned techniques, we derive an efficient policy that can realize sim-to-real generalization.
+In this subsection, we explain our built simulation environment for dynamic object manipulation on conveyor belts.
+Thanks to the appearance noise annealing strategy, the GEM network primarily focuses on the geometric structure of objects, and its reliance on visual appearance is reduced.
+To ensure that the trained network can generalize to diverse unseen objects, we need to generate manipulation demonstrations using 3D assets of varied geometries.
+This is because a key advantage of our method is that while the texture appearances of objects are highly diverse and nearly impossible to exhaustively enumerate, the set of common geometries is relatively limited.
+Therefore, by collecting a set of 3D assets that covers common object geometries, our method can achieve promising generalization.
+This insight empowers us to avoid collecting a very large number of 3D assets, significantly reducing the manual effort required for 3D asset preparation.
+Based on this insight, we select 30 assets with representative geometries from the GSO dataset [10] to generate demonstration data, and these assets are depicted in Fig.
+7 (a).
+We adopt Isaac Gym [31] to build the simulation environment for dynamic object manipulation, as visualized in Fig.
+7 (b).
+We do not employ more visually photorealistic simulators like Isaac Sim [34] because we find their simulation fidelity on object interaction dynamics is inadequate for automatic data generation.
+Hence, in this work, Isaac Sim is mainly utilized to evaluate the out-of-domain generalization capabilities of policies trained using the demonstrations collected in Isaac Gym.
+Since Isaac Gym does not support rendering complex appearance textures, we assign random solid colors to objects during data collection, as illustrated in Fig.
+7 (b).
+By contrast, as depicted in Fig.
+7 (a), objects exhibit realistic textures during the evaluation in Isaac Sim.
+[…]
+To enable automatic demonstration generation without human intervention, we employ CuRobo [45] to generate manipulation trajectories based on the true states of objects obtained from the simulator.
+We collect 2,000 demonstrations for each of the four skills, and every demonstration comprises about ten seconds of manipulation actions, robot proprioception information, and environment observations.
+V Experimental Verifications
+V-A Experimental Settings
+V-A1 Implementation Details
+As described in Section IV-E, we generate 2,000 demonstrations for each skill in Pick, Put, Rotate, and Insert.
+We obtain a total of 8,000 demonstrations for these four skills and utilize them to train networks.
+We optimize the GEM network for 10K optimization iterations.
+The training takes about 14 hours on two RTX4090 GPUs.
+The optimizer is AdamW [30] and the learning rate is 1e41e-4.
+The training batch size is 32.
+The image resolutions of the two RGB-D cameras are both $640\times 480$.
+The action chunk size $N_{a}$, memory token number $l_{m}$, and feature channel number $c$ are set to 20, 200, and 512, respectively.
+GEM controls the robot at a frequency of 20 HZ.
+Since objects are transported successively, this work does not consider scenarios where objects on the conveyor belt are cluttered together.
+Figure 8: The real-robot experiment platform and remote teleoperation handle.
+Specifically, Sub-figure (a) illustrates the real-robot environment for studying dynamic object manipulation.
 GEM controls the robot arm to perform diverse tasks given the observations from the external and wrist RGB-D cameras.
-Sub-figure (b) shows our designed remote handle for teleoperating the robot arm to collect demonstration data.
-IV-A Hardware Setup
-To study the dynamic object manipulation problem, we build a hardware platform as shown in Fig.
-5.
+Sub-figure (b) shows our designed remote handle for teleoperation.
+In the experiments needing real-world tuning data, we utilize this remote handle to collect demonstrations efficiently.
+Figure 9: The 18 objects of diverse geometries selected to conduct real-world experiments.
+All the objects can be used in the evaluation of the skills Pick and Rotate.
+For Put, one of the small-sized objects highlighted in the green box is put in one of the containers in the yellow box.
+The objects in the orange box are for Insert, which only allows two millimeters of insertion error.
+Figure 10: (a) We build a simulation environment in Isaac Sim to evaluate the out-of-domain generalization of policies trained using the demonstrations from Isaac Gym.
+(b) We select 20 additional assets to validate the manipulation performance of policies on unseen objects.
+V-A2 Real-Robot Platform
+To conduct real-robot experiments, we build a hardware platform as shown in Fig.
+8.
+We employ this platform to evaluate the sim-to-real performance of policies trained using simulated demonstrations.
 In this platform, objects are driven by a conveyor belt at an adjustable velocity.
-A robot arm with a gripper is mounted on a side of this conveyor belt to manipulate the target object.
+A robot arm with a gripper is mounted on one side of this conveyor belt to manipulate the target object.
 Two RGB-D cameras are deployed to capture environment observations, one on the robot arm (the wrist camera) and the other on a fixed third-person perspective bracket (the external camera).
@@
 8 (a).
+This handle is used in the experiments needing real-world tuning demonstrations.
 Specifically, we control the motion of the robot arm end-effector based on the relative pose change of the handle derived from the SLAM camera.
@@
 The control signals from the SLAM camera and trigger are wirelessly transmitted to the robot arm by the main control board, thereby enabling flexible remote teleoperation.
-The teleoperation handle is only utilized to collect demonstration data for training the learnable network, which we explain in Section IV-E.
 The handle is not used when testing the method.
-Figure 6: The overall pipeline of GEM.
-GEM primarily consists of two blocks, the object-centric geometry observation and hybrid action control.
-In object-centric geometry observation, we exclude the many factors irrelevant to predicting actions from policy observation, and thus the observation entropy is reduced.
-In hybrid action control, we combine classic visual servo control with learnable network control.
-In the learnable network control, we design the recurrent probabilistic network, which models historical vision observations as memory.
-In this way, the action prediction entropy is decreased.
-Only using hundreds of demonstrations for training, GEM can be generalized to diverse unseen scenes.
-In this figure, the pink modules are implemented based on learnable neural networks and the yellow modules are non-learnable algorithms.
-IV-B GEM Overall Framework
-The overall framework of GEM is illustrated in Fig.
-6.
-It can be observed that GEM comprises two blocks, i.e., object-centric geometry perception and hybrid action control.
-The object-centric geometry perception block extracts 3D point clouds of the target objects in the external and wrist sensor views.
-The design of this block follows the principles summarized in Section III-C, which points out the importance of excluding factors irrelevant to predicting actions from observations.
-In this way, the observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is decreased and less data is required to achieve strong generalization.
-In hybrid action control, we combine a classic visual servo control algorithm [1] and an imitation learning based neural network control algorithm to compute the robot end-effector trajectories.
-Specifically, the visual servo control algorithm is responsible for tracking the target object to manipulate, and the neural network control algorithm (the recurrent probabilistic network) generates manipulation actions of different skills according to input skill type instructions.
-The robot end-effector action is the sum of the actions produced by these two algorithms.
-This design allows the recurrent probabilistic network not need to consider the moving dynamics (like velocity and acceleration) of the target object, further reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-In addition, the developed recurrent probabilistic network can model long-term historical memory, thereby decreasing the action prediction entropy $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$.
-The output actions of the recurrent probabilistic network are represented in Gaussian Mixture Distributions as described in Section III-E.
-Hence, the manipulation multimodal ambiguity problem is addressed.
-To enable GEM to adapt to different robot hand embodiments, the hybrid action control block predicts the translation and orientation of the robot end-effector.
-Based on this prediction, we calculate the corresponding robot joint rotations via inverse kinematics [53], and these rotations are utilized to operate the movements of robot arms.
-This strategy allows for compatibility with various robot embodiments.
-IV-C Object-Centric Geometry Perception
-This subsection describes a perception module that produces observations with low $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-To design this module, we first identify the redundant factors in observations.
-If these factors do not provide necessary information for predicting actions, we can exclude them from observations to reduce $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ while maintaining manipulation performance.
-Fig.
-2 summarizes the primary factors affecting $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-By analyzing them, we first observe that successful manipulation mainly depends on the geometric structures of objects rather than their appearances.
-However, most embodied networks take images as input [46], which conveys visual appearance directly but lacks explicit 3D geometry.
-To handle the missing geometry, we employ uncolored 3D points as policy input.
-These 3D points can be easily accessed via the RGB-D cameras shown in Fig.
-5.
-Besides, some recent deep learning models can produce promising dense depth maps from monocular or binocular images [54, 55], which facilitates obtaining depth.
-In addition, 3D points provide representation less affected by sensor parameters.
-This is because 3D points exist in the real world and are not represented in the pixel coordinate system.
-Since environment background does not affect manipulation interaction, we adopt a lightweight instance segmentation model [56] to segment points of concerned objects and remove background points.
-Such an instance segmentation model is readily obtainable due to the availability of many open-world segmentation models [57].
-Notably, multiple objects could be observed in a single frame.
-For example, in the “pick up the spoon out of the cup” task illustrated in Fig.
-6, the 3D points of both the spoon and cup are important.
-In industrial applications, as manipulation procedures are predetermined, we can define which objects to segment in advance.
-For tasks with unfixed procedures, large language models [26] could be employed to reason about which objects to focus on.
-Notably, a manipulation policy needs to classify the points belonging to different objects, like the cup and spoon in the example shown in Fig.
-6.
-Nevertheless, using uncolored 3D points as policy input makes this classification challenging.
-To address this obstacle, we leverage instance segmentation results to append distinct instance identifiers (IDs) to the points of various objects.
-With the instance IDs, the object classification problem is addressed.
-Summarily, combining the above designs, we propose to use uncolored 3D points of concerned objects appended with corresponding instance IDs as the observation representation with low $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-IV-D Hybrid Action Control
-This subsection explains how to further reduce $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ by devising the hybrid action control block in Fig.
-6.
-The basic idea is to leverage the complementary strengths of classic control and learnable network control.
-Classic control algorithms, such as visual servo control [1], excel at stably and rapidly guiding the robot arm to approach and track moving objects.
-However, they often struggle with collision-rich interactions, such as grasping, inserting, and rotating objects.
-This is because of the complexity of modeling contact dynamics.
-On the other hand, learnable network control methods, particularly those grounded in imitation learning, are adept at handling such interactions by capturing the underlying patterns in expert demonstrations.
-By combining these two approaches, we enable the learnable network to concentrate on generating collision-rich interactions, thereby reducing $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-Figure 7: This figure illustrates how classic control and learnable network control are combined.
-Specifically, classic control is utilized to approach and track the moving target object.
-The learnable network control starts generating manipulation actions once stable tracking is achieved.
-The grippers and target objects marked in gray, dark yellow, and reddish brown correspond to their positions at Moments 1, 2, and 3, respectively.
-The pink gripper indicates the position that would be achieved using classic control alone at Moment 3.
-As described in Section IV-C, the object-centric geometry perception module extracts the 3D points of the target object to manipulate from both the external and wrist sensors.
-Denote the two groups of 3D points as $\{(x_{i}^{e},y_{i}^{e},z_{i}^{e})\}_{i=1}^{N^{e}}$ and $\{(x_{i}^{w},y_{i}^{w},z_{i}^{w})\}_{i=1}^{N^{w}}$, respectively.
-The classic control, consisting of an object state estimator and an end-effector tracking planner as shown in Fig.
-6, operates the robot hand to track the moving target object based on the external view 3D points.
-Specifically, the object state estimator first computes the top centroid of the external view 3D points, denoted as $(x_{c}^{e},y_{c}^{e},z_{c}^{e})$, where $x_{c}^{e}=\frac{1}{N_{e}}\sum\limits_{i=1}^{N_{e}}x_{i}^{e}$, $y_{c}^{e}=\frac{1}{N_{e}}\sum\limits_{i=1}^{N_{e}}y_{i}^{e}$, and $z_{c}^{e}=\max\limits_{i=1}^{N_{e}}z_{i}^{e}$ (the positive z-axis points upward).
-Based on the top centroids at different timestamps, we can estimate the target object velocity $(v_{x},v_{y},v_{z})$ using Gaussian process regression [58] implemented in the object state estimator.
-If the target object is occluded in the external view camera, its real-time top centroid is also predicted by this estimator based on historical centroid records using Gaussian process regression.
-The control target of the end-effector tracking planner is to make the end-effector stably track the motion of the target object.
-In other words, the relative position offset between the end-effector and the target object should be maintained at the preset distances $(x^{p},y^{p},z^{p})$.
-In addition, to ensure the wrist RGB-D camera can clearly view the target object, the orientation of the end-effector should be kept at the preset Euler angles $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$.
-Summarily, the tracking action generated by the end-effector tracking planner comprises three parts, the tracking position $(x_{c}^{e}+x^{p},y_{c}^{e}+y^{p},z_{c}^{e}+z^{p})$, tracking orientation $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$, and tracking velocity $(v_{x},v_{y},v_{z})$.
-Based on the tracking action, the end-effector first approaches the target object, as visualized in Moment 1 of Fig.
-7.
-Then, at Moment 2, the end-effector achieves a stable tracking status.
-In this status, the moving target object appears stationary in the view of the wrist camera.
-After attaining this status, the learnable network (the recurrent probabilistic network) starts to predict manipulation actions for collision-rich interactions with the target object, as shown in Moment 3.
-Since the network generates manipulation actions based on the wrist-view 3D points $\{(x_{i}^{w},y_{i}^{w},z_{i}^{w})\}_{i=1}^{N^{w}}$, the original dynamic object manipulation problem is transformed into a static object manipulation problem in the view of the network.
-Hence, the motion pattern of the target object is excluded from the factors affecting observations, meaning $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is reduced.
-Besides, this design fixes the relative position between the end-effector and the target object in the beginning of manipulation action generation as $(x^{p},y^{p},z^{p})$, which further decreases $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-The manipulation action predicted by the network includes the position offset $(x^{w},y^{w},z^{w})$ and orientation offset $(\theta^{w}_{x},\theta^{w}_{y},\theta^{w}_{z})$.
-The end-effector control target output by the hybrid action control block is the sum of the tracking action and manipulation action.
-Therefore, the end-effector control target consists of the position target $(x_{c}^{e}+x^{p}+x^{w},y_{c}^{e}+y^{p}+y^{w},z_{c}^{e}+z^{p}+z^{w})$, orientation target $(\theta^{p}_{x}+\theta^{w}_{x},\theta^{p}_{y}+\theta^{w}_{y},\theta^{p}_{z}+\theta^{w}_{z})$, and velocity target $(v_{x},v_{y},v_{z})$.
-Afterwards, the end-effector control target is converted into robot joint rotations based on inverse kinematics.
-This design enables a single GEM system to be applied to different robot embodiments.
-Figure 8: The overall framework of the proposed recurrent probabilistic network, which takes the 3D points of concerned objects as input and includes a memory encoder to manage the historical observation memory.
-IV-E Recurrent Probabilistic Network
-Section IV-D has described how classic control and learnable network control are combined in the hybrid motion control block.
-This part further elaborates the structure details of the recurrent probabilistic network used in learnable network control.
-The overall framework of this network is illustrated in Fig.
-8.
-To concentrate on the geometry of objects rather than appearance, the recurrent probabilistic network takes the wrist view 3D points produced by the object-centric geometry perception block as the main input.
-For the 3D points perceived at timestamp $t$, it first employs a point feature extractor to encode the 3D points into semantic-rich feature tokens $F_{t}\in\mathbb{R}^{l_{p}\times c}$, where $l_{p}$ denotes the token number and $c$ is the token length.
-In this work, we model the point feature extractor with Sonata [25], a recent promising point network.
-However, the input of Sonata should be colored points, while ours are uncolored points appended with instance IDs.
-To bridge this gap, we feed the instance IDs into Sonata as surrogate color information.
-Therefore, the points of various object instances in Fig.
-8 are visualized with different colors.
-The memory encoder shown in Fig.
-8 is to fuse $F_{t}$ and the memory feature at timestamp t1t-1, denoted as $M_{t-1}\in\mathbb{R}^{l_{m}\times c}$, into the new memory feature $M_{t}\in\mathbb{R}^{l_{m}\times c}$.
-To this end, the memory encoder first employs a Transformer decoder [59] to conduct feature interaction between $M_{t-1}$ and $F_{t}$.
-Specifically, $M_{t-1}$ is used as the query embedding, and $F_{t}$ is the key and value embeddings.
-The feature output by this Transformer decoder is transformed into the forget score $G_{t}\in\mathbb{R}^{l_{m}\times 1}$ and hidden feature $H_{t}\in\mathbb{R}^{l_{m}\times c}$ by two multilayer perceptron layers (MLPs).
-Afterwards, in the Mix module visualized in Fig.
-8, $M_{t}$ is calculated following:
-$M_{t}=\sigma(G_{t})\odot H_{t}+(1-\sigma(G_{t}))\odot M_{t-1},$ (7)
-where $\sigma(\cdot)$ represents the Sigmoid function.
-At the beginning of manipulation, the memory feature is initialized as a zero embedding.
-Taking these designs, the historical memory is updated in an auto-regressive style, and $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is decreased.
-Then, $M_{t}$, $F_{t}$, and a skill type embedding $S_{t}\in\mathbb{R}^{1\times c}$ are concatenated and input to three Transformer decoders.
-$S_{t}$ is a learnable token specifying which skill to perform, and the supported skills include Pick, Place, Rotate, and Insert.
-The feature interaction process in decoders is similar to the implementation in ACT [8], where a chunk of $N_{a}$ action tokens is utilized to conduct feature interaction with the input feature.
-Then, the action tokens are decoded as the action probability distributions of future $N_{a}$ timestamps and a binary finish flag.
-As described in Section III-E, the action probability distributions are represented as the means, variances, and component weights of Gaussian mixture distribution.
-The finish flag indicates whether the execution of the manipulation skill has been completed.
-If completed, the network stops generating actions until the next time of performing a skill.
-In training, the network is updated following Eq.
-(3) as described in Section III-E.
-During testing, the action with the maximum probability is selected for execution.
-Figure 9: This figure illustrates our simulated dynamic object manipulation environment.
-(a) An overview of the environment, where diverse objects are moving on a conveyor belt.
-(b) We collect 3D assets in 30 representative geometric structures to collect demonstrations.
-(c) The Pick skill.
-(d) The Put skill.
-(e) The Rotate skill.
-(f) The Insert skill.
-IV-F Efficient Demonstration Collection
-Combining the aforementioned techniques, we derive a low-entropy manipulation system, which exhibits stronger generalization with less demonstration data.
-In this part, we introduce how the demonstration data is efficiently collected to support various manipulation skills within a single system.
-Demonstration data can be collected in both simulated environments and the real world.
-Data collection in simulated environments is usually more efficient, but the visual fidelity of existing simulators is still limited.
-This limitation restricts existing simulated data generation.
-Thankfully, our network takes uncolored 3D points as input, avoiding the visual fidelity problem.
-Therefore, our method can generalize to real-world applications even though only using simulated data.
-Nevertheless, there is no public simulated environment for dynamic object manipulation.
-To bridge this gap, we build the first such environment based on Isaac Gym [32].
-As shown in Fig.
-9 (a), the environment layout is similar to the real platform in Fig.
-5.
-As mentioned in Section IV-C, geometric structure information of objects is essential for successful manipulation.
-To enable our method to generalize to objects in diverse geometries, we gather 3D assets of 30 representative geometric structures from the GSO dataset [60] to produce demonstration data.
-These 3D assets are visualized in Fig.
-9 (b).
-This work supports the four common skills including Pick, Place, Rotate, and Insert.
-These four skills are shown in Fig.
-9 (c)-(f), respectively.
-Specifically, Pick requires the robot to pick up the moving object from the conveyor belt.
-Place needs the robot to put an object into a moving container on the conveyor belt.
-Rotate demands the robot to rotate the moving object to a target orientation.
-The target orientation in this work is prompted to the network by providing a reference observation, which is the point cloud of the manipulated object in the goal 3D orientation from the perspective of the network.
-Insert involves inserting an object into a container with a precisely sized opening, ensuring the object is aligned at the specified pose.
-The Insert task illustrated in Fig.
-9 (f) only allows two millimeters of position error.
-During gathering demonstrations, all the 30 assets can be used to collect Pick data.
-For Put, the 12 small-sized assets highlighted by a green box in Fig.
-9 (b) are put into the 5 huge-sized containers marked by a yellow box.
-In Rotate, the 6 assets lacking rotational symmetry are utilized to collect data.
-For Insert, due to the size matching requirement between the insert object and the container, we only utilize the assets shown in Fig.
-9 (f).
-Notably, the 3D assets are colored with random pure colors instead of being rendered with more realistic appearances.
-This is because our method takes uncolored point clouds as input, so there is no need to spend computational resources on rendering appearances.
-Among these four skills, Pick requires human knowledge to decide reasonable grasp positions.
-For example, in Fig.
-9 (c), the robot should grasp the hammer handle rather than the hammer head.
-Due to this requirement, the Pick data should be collected through human teleoperation.
-Specifically, we use our designed teleoperation handle shown in Fig.
-5 (b) to collect 20 demonstrations for every 3D asset.
-In this way, 600 Pick demonstrations are obtained.
-Different from Pick, the demonstration data of the other three skills can be generated automatically through preprogrammed manipulation procedures.
-This is because we have the true state information of all objects in simulated environments.
-Hence, we build an automatic pipeline to generate data for these three skills without human intervention.
-A total of 2000 demonstrations are derived for each of these three skills.
-V Experimental Verifications
-V-A Experimental Settings
-V-A1 Implementation Details
-We train the recurrent probabilistic network using the 6,600 demonstrations collected in Section IV-F for 10K optimization iterations.
-The training takes about 14 hours on a single RTX4090 GPU.
-The optimizer is AdamW [61] and the learning rate is 1e41e-4.
-The training batch size is 32.
-The image resolutions of the two RGB-D cameras are both $640\times 480$.
-The action chunk size $N_{a}$, memory token number $l_{m}$, and feature channel number $c$ are set to 20, 200, and 512, respectively.
-GEM controls the robot at a frequency of 20 HZ.
-V-A2 Testing Protocols
-We test GEM in both the simulated environment in Fig.
-9 and the real platform in Fig.
-5.
+To evaluate the capabilities of GEM on manipulating objects in various geometries, we select 18 objects as shown in Fig.
+9 to conduct experiments.
+All the objects can be used to validate the skills Pick and Rotate.
+For Put, the robot needs to put one of the small-sized objects highlighted in the green box into one of the containers marked in the yellow box.
+The objects in the orange box are for Insert, which are in various sizes and only allow two millimeters of insertion error.
+V-A3 Simulation Platform
+Although real-robot experiments directly reflect robot performance in the real world, it is inconvenient to adjust environment variables to conduct an ablation study.
+In contrast, simulation experiments allow controlling environmental variables precisely, so we also evaluate our method in simulation.
+As illustrated in Fig.
+7 (a) and (b), we gather objects of diverse geometries and build a simulation environment in Isaac Gym to generate training demonstrations.
+We can also use this environment to validate the in-domain performances of trained policies.
+To further test the out-of-domain generalization results, we develop a simulation environment in Isaac Sim, as depicted in Fig.
+10 (a).
+We can find that the visual fidelity of the rendered environment in Isaac Sim is promising, ensuring the behaviors of the evaluated policies are similar to those in the real world.
+Notably, a problem with using Isaac Sim is the occasional occurrence of abnormal collision dynamics.
+For example, a gripper might lightly touch a cup, causing it to jump or move unrealistically.
+Due to this problem, we do not choose Isaac Sim to build the automatic demonstration generation pipeline.
+During evaluating policies in Isaac Sim, we manually inspect such simulation anomalies and exclude abnormal cases from the experimental statistics.
+In addition, as shown in Fig.
+10 (b), we select 20 new objects to validate the manipulation success rates of policies on unseen objects.
+V-A4 Testing Protocols
+We analyze policies using both the aforementioned simulated environments and a real-world platform.
 The primary evaluation metric is the task completion success rate.
@@
 In Put and Insert, an object and a container are moving on a conveyor belt.
-Success is defined as the robot arm placing the object inside the container.
+Success is defined as the robot arm places the object inside the container.
 For Rotate, the robot hand grasps and rotates the object on the conveyor belt to the target orientation.
@@
 The default moving speed of the conveyor belt is 0.1 m/s.
-To mitigate the influence of random errors, we conduct each experiment for 100 times to compute the average success rate.
-Figure 10: This figure illustrates the 14 selected objects in diverse geometric structures for out-of-domain manipulation evaluation.
-All the assets are used for Pick.
-In Put, the small-sized assets highlighted in a green box are required to be put in the containers marked in yellow.
-The ten 3D assets lack of rotational symmetry are employed in Rotate.
-Figure 11: The 18 objects in diverse geometries that are used to conduct the real-world manipulation experiments.
-The left 9 objects can be used for collect demonstrations and are used in in-domain evaluation.
-The right 9 objects are for out-of-domain validation.
-All the objects can be used in Pick.
-For Put, one of the small-sized objects highlighted in the green box is put in one of the containers marked in the yellow box.
-The USB adapter and spoon are employed in Rotate.
-The objects in the orange box are for Insert, which also only allows two millimeters of manipulation error.
+To mitigate the influence of random errors, we repeat each experiment for 100 times to compute the average success rate.
+In simulation, we validate policies based on the in-domain and out-of-domain settings.
+In the in-domain setting, the policies are evaluated in Isaac Gym to manipulate the objects that have appeared in training demonstrations.
+By contrast, for the out-of-domain setting, the policies manipulate the unseen objects shown as Fig.
+10 (b) in Isaac Sim environments.
+For real-robot experiments, we test policies with the tuned sim-to-real and zero-shot sim-to-real settings.
+In tuned sim-to-real, the policies pre-trained with simulated data are further fine-tuned with 10 real-world demonstrations before evaluation.
+Differently, the zero-shot sim-to-real setting requires the policies to operate a real robot without real-world data tuning.
 V-B Method Comparison
@@
 Although the classic kinematics-based methods can also handle dynamic object manipulation after precise system identification and careful parameter tuning, they demand special algorithm designs for different tasks and testing environments.
-This task-dependent characteristic violates the focus of this work, how to improve the generalization of a manipulation system.
+This task-dependent characteristic violates the focus of this work, how to build a policy that is generalizable across various scenarios.
 Therefore, we mainly compare GEM with the recent learning-based methods, which present better generalization.
-The compared methods include AnyGrasp [37], ConvMLP[62], RVT-2 [63], ACT [8], Diffusion Policy (DP) [46], OpenVLA [11], and PI0 [12].
+The compared methods include AnyGrasp [12], ConvMLP[57], RVT-2 [14], ACT [60], Diffusion Policy (DP) [7], Diffusion Policy 3D (DP3) [56], OpenVLA [24], and PI0 [4].
 Among them, AnyGrasp can generate grasp pose candidates based on the 3D points of the target object.
 Then, the robot arm selects the grasp pose with the highest confidence and grasps the target based on kinematic control.
-Since AnyGrasp only predicts the grasp pose of each object for once, the grasp is executed in an open-loop manner.
+Since AnyGrasp only predicts the grasp pose of each object once, the grasp is executed in an open-loop manner.
 RVT-2 is a representative two-stage method that takes RGB-D information as input and processes this information based on point cloud reconstruction.
+Similar to the GEM network, DP3 takes 3D points as environmental observation, but it does not employ visual color feature.
 The inputs of the other compared methods are the wrist and external view images.
@@
 […]
-AnyGrasp 0.41 - - - - 0.43 - - - - - 4.35
-ConvMLP 0.06 0.19 0.05 0.00 0.08 0.00 0.08 0.00 0.00 0.02 0.05 17.54
-ConvMLP* 0.28 0.48 0.19 0.01 0.24 0.06 0.35 0.02 0.00 0.11 0.18 13.16
-RVT-2 0.08 0.24 0.03 0.00 0.09 0.00 0.14 0.00 0.00 0.04 0.07 31.48
-RVT-2* 0.47 0.62 0.15 0.27 0.38 0.18 0.50 0.00 0.12 0.20 0.29 19.69
-ACT 0.14 0.29 0.10 0.00 0.13 0.05 0.22 0.00 0.00 0.07 0.10 43.48
-ACT* 0.62 0.75 0.53 0.44 0.59 0.26 0.65 0.08 0.27 0.32 0.46 23.70
-DP 0.09 0.24 0.07 0.00 0.10 0.00 0.17 0.00 0.00 0.04 0.07 27.32
-DP* 0.55 0.70 0.46 0.36 0.52 0.18 0.54 0.05 0.23 0.25 0.39 17.86
-OpenVLA 0.11 0.26 0.08 0.00 0.11 0.07 0.24 0.01 0.00 0.08 0.10 1.61
-OpenVLA* 0.64 0.74 0.46 0.35 0.55 0.35 0.68 0.06 0.29 0.35 0.45 1.58
-PI0 0.23 0.36 0.18 0.00 0.19 0.12 0.29 0.10 0.00 0.13 0.16 1.97
-PI0* 0.68 0.81 0.59 0.65 0.68 0.50 0.75 0.14 0.43 0.46 0.57 1.92
-GEM (Ours) 0.88 0.98 0.83 0.80 0.87 0.77 0.91 0.70 0.72 0.78 0.83 23.81
-TABLE II: Success rate with previous methods in real world tasks.
+AnyGrasp* 0.41 - - - - 0.46 - - - - - 4.35
+ConvMLP 0.06 0.19 0.05 0.00 0.08 0.00 0.03 0.00 0.00 0.01 0.05 17.54
+ConvMLP* 0.28 0.48 0.19 0.01 0.24 0.03 0.24 0.02 0.00 0.08 0.16 13.16
+RVT-2 0.08 0.24 0.03 0.00 0.09 0.00 0.07 0.00 0.00 0.02 0.05 31.48
+RVT-2* 0.47 0.62 0.15 0.27 0.38 0.12 0.40 0.00 0.05 0.14 0.26 19.69
+ACT 0.14 0.29 0.10 0.00 0.13 0.02 0.19 0.00 0.00 0.05 0.09 43.48
+ACT* 0.62 0.75 0.53 0.44 0.59 0.20 0.39 0.07 0.23 0.22 0.41 23.70
+DP 0.09 0.24 0.07 0.00 0.10 0.00 0.15 0.00 0.00 0.04 0.07 27.32
+DP* 0.55 0.70 0.46 0.36 0.52 0.17 0.32 0.03 0.18 0.18 0.35 17.86
+DP3 0.31 0.50 0.29 0.00 0.28 0.19 0.40 0.22 0.00 0.20 0.24 29.40
+DP3* 0.72 0.85 0.77 0.70 0.76 0.57 0.74 0.62 0.44 0.59 0.68 18.87
+OpenVLA 0.11 0.26 0.08 0.00 0.11 0.06 0.18 0.01 0.00 0.06 0.09 1.61
+OpenVLA* 0.64 0.74 0.46 0.35 0.55 0.31 0.64 0.05 0.24 0.31 0.43 1.58
+PI0 0.23 0.36 0.18 0.00 0.19 0.09 0.24 0.06 0.00 0.10 0.15 1.97
+PI0* 0.68 0.81 0.59 0.65 0.68 0.47 0.70 0.11 0.38 0.42 0.55 1.92
+GEM (Ours) 0.94 1.00 0.90 0.86 0.93 0.86 0.96 0.90 0.83 0.89 0.91 23.81
+TABLE II: Success rate comparison with previous methods in real world tasks.
 […]
-AnyGrasp 0.12 - - - -
-ConvMLP 0.00 0.00 0.00 0.00 0.00
-ConvMLP* 0.06 0.38 0.02 0.00 0.12
-RVT-2 0.02 0.15 0.00 0.00 0.04
-RVT-2* 0.17 0.59 0.08 0.04 0.22
-ACT 0.08 0.23 0.05 0.00 0.09
-ACT* 0.53 0.76 0.40 0.35 0.51
-DP 0.05 0.17 0.01 0.00 0.06
-DP* 0.46 0.70 0.32 0.25 0.43
-OpenVLA 0.00 0.00 0.00 0.00 0.00
-OpenVLA* 0.29 0.35 0.14 0.00 0.20
-PI0 0.00 0.00 0.00 0.00 0.00
-PI0* 0.37 0.58 0.21 0.00 0.29
-GEM (Ours) 0.92 0.96 0.79 0.74 0.85
+AnyGrasp* 0.45 - - - -
+ConvMLP 0.08 0.25 0.07 0.00 0.10
+ConvMLP* 0.36 0.55 0.28 0.13 0.33
+RVT-2 0.13 0.30 0.09 0.00 0.13
+RVT-2* 0.51 0.68 0.41 0.28 0.47
+ACT 0.18 0.34 0.20 0.00 0.18
+ACT* 0.65 0.76 0.55 0.49 0.61
+DP 0.15 0.29 0.13 0.00 0.14
+DP* 0.61 0.69 0.56 0.32 0.55
+DP3 0.36 0.45 0.39 0.04 0.31
+DP3* 0.70 0.80 0.72 0.61 0.71
+OpenVLA 0.00 0.02 0.05 0.00 0.02
+OpenVLA* 0.32 0.39 0.20 0.03 0.24
+PI0 0.03 0.08 0.14 0.00 0.06
+PI0* 0.40 0.48 0.56 0.11 0.39
+GEM (Ours) 0.98 1.00 0.93 0.89 0.95
 V-B1 Simulated Comparison
-This part compares GEM with the aforementioned counterparts in the simulation tasks described in Section V-A.
+This part compares GEM with the aforementioned counterparts in simulation.
 Since AnyGrasp is for generating grasp poses, we only test it in Pick tasks.
-The algorithms in the hybrid action control module are utilized to help AnyGrasp handle the moving dynamics of objects.
-The other compared methods are trained using the same demonstrations as our recurrent probabilistic network.
-Because the compared models have no designs for dynamic object manipulation, we test them in two modes, the sole mode and hybrid mode.
-The sole mode means the robot movement is solely controlled by a network output.
-In the hybrid mode, the network output is combined with visual servo tracking like our developed hybrid action control module.
-In this way, the original dynamic object manipulation problem is converted into a static problem, which favors the performances of the compared methods.
-As mentioned in Section IV-F, we collect 30 diverse assets to generate demonstrations and train the compared methods except AnyGrasp.
-To verify their out-of-domain generalization on unseen geometries, we additionally select 14 new assets, as shown in Fig.
-10, to conduct experiments.
-The out-of-domain Insert validation is conducted using blocks similar to the ones shown in Fig.
-9 but we randomly change the block sizes.
+The other compared methods are trained using the same simulated demonstrations as our GEM network.
+Since the compared models have no designs for dynamic object manipulation, we test them in two modes, the sole mode and hybrid mode.
+The sole mode means the robot movement is solely controlled by their network outputs.
+In the hybrid mode, we utilize a segmentation model [21] to estimate the spatial position of the target object and then employ a visual servo control algorithm [38] that controls the end-effector to track the target object.
+The control signal is the sum of the network output and the visual servo control module, which shares a similar process as our proposed manipulation action decomposition strategy.
+In this way, the original dynamic object manipulation problem is converted into a static object manipulation problem, significantly improving the performance of the compared methods.
+As AnyGrasp is an open-loop method, we only report its results in the hybrid mode.
+As described in Section V-A, we compare these policies in the in-domain setting (manipulate seen objects in Isaac Gym) and out-of-domain setting (manipulate unseen objects in Isaac Sim).
 The success rates and inference speeds of all the methods are presented in Table I.
-Specifically, the $2_{\rm nd}-6_{\rm th}$ columns are the success rates of in-domain evaluations based on the assets shown in Fig.
-9 (b).
-The $7_{\rm th}-11_{\rm th}$ columns correspond to the out-of-domain success rates of using the assets in Fig.
-10.
+Specifically, the $2_{\rm nd}-6_{\rm th}$ columns are the success rates of the in-domain setting.
+The $7_{\rm th}-11_{\rm th}$ columns correspond to the out-of-domain setting.
 Avg_in, Avg_out, and Avg refer to the average success rates of in-domain tests, out-of-domain tests, and all tests, respectively.
 The inference speeds reported in the $13_{\rm th}$ column are the model inference frequencies tested on an RTX4090 GPU.
-An asterisk (*) following the method name indicates that the method is validated in the hybrid mode.
+An asterisk (*) following a method name indicates that the method is validated in the hybrid mode.
 The hybrid mode increases the inference time compared to the sole mode (denoted without the asterisk) because of the additional adoption of an instance segmentation model.
-According to the results, we can find that GEM achieves the best success rates in performing all the skills.
-This is partly because of the efficient structure of the recurrent probabilistic network on perceiving object geometries based on points.
-In addition, it is observed that the methods in the hybrid mode significantly outperform the ones in the sole mode, which is attributed to that our hybrid action control module alleviates the manipulation challenge.
+According to the results, we can find that GEM achieves the best results in performing all the skills.
+This is partly because of the design that helps the GEM network prioritize geometry information.
+In addition, it is observed that the methods tested in the hybrid mode significantly outperform those of the sole mode, which suggests the effectiveness of our action decomposition strategy.
 For AnyGrasp, we find it generates promising pick poses.
-However, it works in an open-loop manner, which restricts its effectiveness.
+However, it works in an open-loop manner, which restricts its real-time response to object pose changes.
 Though RVT-2 uses depth information like our method, its network architecture is inefficient in geometry information exploitation, resulting in its unsatisfactory success rates.
-Comparing the results of the VLA models and other lightweight networks like ACT, we observe VLAs generally present better out-of-domain generalization thanks to they are pre-trained on numerous data.
+Comparing the results of the VLA models and other lightweight networks like ACT, we observe that VLAs generally present better out-of-domain generalization because they are pre-trained on numerous data.
 However, the speeds of VLAs are slower.
-The slow inference does not affect the success rates significantly in simulated environments because the simulation is paused during model inference, but it harms performances in real tasks significantly.
+The slow inference does not affect the success rates significantly in simulated environments because the simulation is paused during model inference, but it harms performance in real-world tasks significantly.
+Compared with the other policies, the success rates of DP3 are promising although its structure is simple.
+This is because it takes 3D points without visual texture as input, and this strategy also prioritizes 3D geometry.
 V-B2 Real Robot Comparison
-This study compares GEM with the aforementioned counterparts in real-world tasks using the experiment platform introduced in Fig 5.
-The objects illustrated in Fig.
-11, which are in diverse geometries and sizes, are selected to conduct this comparison.
-Specifically, we use the left nine objects in Fig.
-11 to collect 20 demonstrations for every skill.
-Then, the total of 80 demonstrations are utilized to fine-tune the policies that have been trained using the simulation data.
-Finally, employing the same nine objects, we test their in-domain performances in the hybrid mode.
-We do not compare the out-of-domain results in this experiment because we find the generalization of the compared methods is generally limited.
-This is because the color differences between in-domain and out-of-domain objects disturb their action prediction.
-In the simulation, all the objects are rendered with pure colors, so the compared methods generalize successfully to some extent according to the results in Table I.
-Nevertheless, in real-world validation, the appearances between in-domain and out-of-domain objects are quite different.
-Differently, GEM can directly generalize to real world-tasks even though trained utilizing only simulated data, according to the sim-to-real results in Section V-E.
-This result highlights the advantage of using uncolored 3D points.
-The real robot comparison result is presented in Table II, where we can find that GEM presents high success rates after tuning using only 80 demonstrations.
-Besides, the results of the compared methods in the hybrid mode significantly outperform the results in the sole mode, indicating the value of combining classic control and learning algorithms.
+This study compares GEM with the aforementioned counterparts in real-robot tasks using the platform shown in Fig.
+8 (a).
+We only report the results in the tuned sim-to-real setting, which employs 10 real-world demonstrations to fine-tune policies before evaluation.
+This is because we find the compared methods mostly cannot generalize to real-world manipulation after training with simulated data, though GEM generalizes well since it prioritizes geometry information.
+The tuning step adapts the policy to the real-world embodiment and environment observations.
+Similar to the simulation experiments, we validate the compared methods in both the sole and hybrid modes.
+The real robot comparison results are presented in Table II, where we can find that GEM presents high success rates.
+According to the results, the sim-to-real performance of GEM outperforms the compared counterparts by large margins.
+Besides, the results of the compared methods in the hybrid mode significantly outperform the results in the sole mode, indicating the value of our manipulation action decomposition.
 Moreover, we can find that although PI0 behaves better than ACT in simulated environments, its success rates are lower in real robot experiments.
@@
 Therefore, fast inference speed is important for models used in dynamic object manipulation.
-Figure 12: Subfigure (a) illustrates the MazeNav task, where a robot is initialized in a random grid of the $5\times 5$ maze and needs to find a path to reach the yellow star.
-The yellow star is always in the botton-right corner of the maze.
-Subfigure (b) visualizes the action of the expert policy in every grid.
-V-C Theory Verification
-This subsection validates the entropy analysis framework described in Section III, i.e., how the changes of observation entropy $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and action prediction entropy $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ affect the target distribution approximation precision.
-As the entropy of robotic manipulation is too complex to quantify, we design a toy task named MazeNav to study this problem.
-Specifically, as shown in Fig.
-12 (a), a robot is initialized in a random grid of the $5\times 5$ maze and needs to find a path to reach the destination marked by a yellow star.
-The destination is always in the bottom-right corner of the maze.
-The action space includes moving up, down, left, and right.
-The robot is not allowed to move beyond the maze boundaries.
-As illustrated in Fig.
-12 (b), a deterministic expert policy is built to generate demonstrations.
-Based on the demonstrations, a student policy is to estimate the expert policy action distribution by calculating the frequency of each action taken in each state observation.
-If an action is never performed given a state, the student policy assigns equal possibilities to all candidate actions in this observation state.
-Besides the robot position, the state observation of the student policy includes an integer variable $n_{m}$ randomly sampled from $1,N_{m}]$.
-As a random variable, $n_{m}$ is irrelevant to action prediction.
-We can change $N_{m}$ to adjust $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-To alter $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, we introduce a threshold variable $\eta\in(0,1)$.
-During collecting demonstrations, if a variable sampled between $(0,1)$ is smaller than $\eta$, the action is randomly selected from the action space rather than decided by the expert policy.
-Hence, by increasing $\eta$, $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is enhanced.
-Figure 13: This figure illustrates the relation between $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and the target distribution approximation precision.
-$\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is adjusted by changing $N_{m}$.
-The approximation precision is measured by the KL divergence between the expert distribution and student distribution.
-Figure 14: This figure illustrates the relation between $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ and the target distribution approximation precision.
-$\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is adjusted by changing $\eta$.
-The approximation precision is measured by the KL divergence between the expert distribution and student distribution.
-V-C1 Observation Entropy Verification
-By adjusting $N_{m}$ and the demonstration trajectory number, this part studies the relation between $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ and the expert distribution approximation precision.
-Specifically, $\eta$ is fixed as 0, meaning that there is no action prediction ambiguity.
-To study the approximation performance under various $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ values, we repeat this experiment by setting $N_{m}$ to 1, 3, and 5 separately.
-In each experiment, the expert policy generates 10, 20, 30, 40, and 50 demonstration trajectories to update the student policy, respectively.
-After the update, the expert distribution approximation precision is measured by the KL divergence between the expert and student action distributions.
-To eliminate the influence of random factors, we set different seeds and repeat each experiment five times to calculate the standard deviation.
-The obtained KL divergence values and standard deviations under different $N_{m}$ and demonstration trajectory numbers are visualized in Fig.
-13.
-We can observe that increasing the demonstration number improves the approximation performance and is helpful in reducing the standard deviation.
-In addition, by increasing $N_{m}$, which means larger $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$, the approximation precision deteriorates when using the same number of demonstrations.
-These observations are consistent with the theoretical analysis in Section III-C.
-V-C2 Action Prediction Entropy Verification
-This part explores the relation between $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ and the expert distribution approximation precision.
-Specifically, $N_{m}$ is fixed as 1, indicating that the policy observation does not include redundant information.
-We set $\eta$ to 0.0, 0.3, and 0.6 separately to adjust $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$.
-Similar to the previous experiment, we update the policy using 10, 20, 30, 40, and 50 demonstration trajectories, respectively.
-The experiment of each hyper-parameter is repeated five times to compute the mean KL divergence and standard deviation.
-The experimental results of different $\eta$ and demonstration numbers are shown in Fig.
-14.
-According to the results, we can find that a larger $\eta$, which indicates bigger $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$, leads to poorer distribution estimation precision.
-In addition, by incorporating more demonstrations, the student policy can still estimate a promising expert policy distribution even though $\eta$ is high.
-These results confirm the theoretical analysis in Section III-D.
-Figure 15: This experiment suggests that the student policy can perfectly estimate the action choices of the expert policy if sufficient demonstrations are provided, even though the demonstrations are very noisy (high $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$) and contain much irrelevant observation (high $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$).
-Specifically, the subfigure (a) visualizes the expert policy.
-Subfigures (b)-(d) show the student policies updated with 50, 200, and 1000 demonstrations, respectively.
-The student policy using 1000 demonstrations perfectly replicates the expert policy.
-V-C3 Effect of Data Scaling Up
-This experiment aims to validate the hypothesis in Section III-B that the optimal actions can be precisely approximated by scaling up expert demonstrations.
-To study this hypothesis, we set $N_{m}$ to 10, which leads to high $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
-Besides, $\eta$ is 0.9, meaning the robot performs random actions in 90% of the steps.
-This results in a large $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$.
-We update the student policy with 50, 200, and 1000 demonstrations.
-Then, we visualize the student policy actions corresponding to the biggest probabilities under various states in Fig.
-15 (b)-(d).
-We can find that even though the demonstration data is highly noisy, the student policy still perfectly replicates the expert policy using 1000 demonstrations.
-This result confirms the hypothesis that the optimal actions can be precisely approximated with sufficient demonstrations.
-TABLE III: Comparison between point and imgae observations.
+TABLE III: Analysis on the appearance noise annealing
 […]
-Image $\times$ 0.75 0.60 0.64 0.33
-Image $\checkmark$ 0.41 0.28 0.38 0.14
-Point $\times$ 0.79 0.71 0.73 0.66
-Point $\checkmark$ 0.87 0.78 0.85 0.74
-Figure 16: This figure visualizes some instance segmentation cases during manipulation, where we can find that the segmentation masks are precise.
-V-D Component Analysis
-This subsection ablates the various designs in GEM.
-Due to the space limit, when simultaneously reporting the in-domain and out-of-domain results in simulated and real environments, we report the average success rates of all skills.
-V-D1 Analysis on Point based Representation
-Different from previous works commonly taking 2D images as input, our method utilizes 3D points.
-This design is based on the intuition that accurate manipulation relies on the geometry of objects, and 3D points better deliver such information.
-This experiment is to validate this intuition.
-The results are presented in Table III, where the first two rows of results correspond to the models based on images, and the last two rows are the models using points.
-When the input is in the image modality, the feature extractor of the recurrent probabilistic network is changed to DINOv2 [64], a widely adopted network due to its discriminative representation.
-The second column indicates whether the background regions are removed based on instance segmentation results.
-“Sim_in”, “Sim_out”, “Real_in”, and “Real_out” refer to the average success rates in the settings of in-domain simulation, out-of-domain simulation, in-domain real world, out-of-domain real world, respectively.
-According to the results, we can first find that the models taking 3D points as input outperform the ones with images, which confirms our intuition.
-Secondly, it is observed that the operation of removing background regions favors point based models but harms the performance of image based models.
-This is because when the input is an image, the model demands background texture to infer the depth of the target object.
-By contrast, the model does not need background since the object position has been known from 3D points.
-V-D2 Analysis on Instance Segmentation
-The results in Table III have revealed how the operation of removing background based on instance segmentation affects manipulation success rates.
-This part analyzes the efficiency and robustness of our employed instance segmentation model, YOLOv11.
-Since the official YOLOv11 only supports segmenting objects in the categories defined in the COCO dataset [65], we need to collect a few additional instance segmentation data to enable YOLOv11 to handle our concerned categories.
-Thankfully, the development of SAM2 [57] makes this data collection process efficient.
-Specifically, based on SAM2, we only need to record a video for every concerned category of objects and click a point on each object to generate segmentation masks.
-In this way, obtaining tuning data totally takes several minutes.
-The inference speed of YOLOv11 evaluated on an RTX4090 GPU is more than 100FPS, meeting the real-time requirement.
-Afterwards, to evaluate the segmentation precision of the tuned YOLOv11, we collect a dataset consisting of 1800 images for the objects shown in Fig.
-11 (200 frames per object).
-Then, we utilize this dataset to validate the segmentation precision of YOLOv11.
-Following the testing protocol in COCO, YOLOv11 achieves a Mask AP of 94.5% in this dataset, suggesting that the segmentation precision YOLOv11 satisfies the manipulation task demand.
-Furthermore, we visualize some challenging cases in Fig.
-16 during manipulation.
-It is found that the segmentation quality is promising.
-Figure 17: This figure visualizes the tracking dynamics of the end-effector under different target velocities.
-We can find that the end-effector stably tracks objects moving at 0.2 m/s but fails to keep up with objects moving at 0.6 m/s.
-TABLE IV: Manipulation performance in various moving speeds.
+GEM-N 0.91 (0.87) 0.70 (0.24) 0.86 (0.77) 0.65 (0.16)
+GEM-S 0.86 (0.69) 0.84 (0.63) 0.91 (0.74) 0.83 (0.60)
+GEM-O 0.93 (0.86) 0.89 (0.83) 0.95 (0.89) 0.87 (0.72)
+V-C Component Analysis
+This subsection ablates the designs in GEM.
+When reporting the results, the four evaluation settings described in Section V-A, including simulated in-domain, simulated out-of-domain, real-world tuned sim-to-real, and real-world zero-shot sim-to-real, are denoted as Sim_in, Sim_out, Real_tune, and Real_zero, respectively.
+Due to the space limit, when reporting the manipulation success rates of these four settings, we report the average success rates of all skills.
+V-C1 Analysis on Appearance Noise Annealing
+This part analyzes how the appearance noise annealing strategy improves the sim-to-real generalization of GEM.
+As described in Section IV-B, we apply strong color perturbation to 3D points at the beginning of training and gradually remove the noise as training progresses.
+To study this strategy, we compare its performance against two baselines, one trained without any color perturbation and another trained with strong color perturbation applied throughout the entire training process.
+The experimental results are presented in Table III.
+The original GEM implementation, the GEM without color perturbation, and the GEM with strong color perturbation all the time are denoted as GEM-O, GEM-N, and GEM-S, respectively.
+The first number in every table cell is the average success rate of four skills, and the second number in a bracket represents the success rate of the Insert skill.
+We separately list the success rates of the Insert skill because this skill requires highly precise operation and the numerical result change is obvious.
+Comparing GEM-O and GEM-N, we can observe that the appearance noise annealing strategy boosts the generalization of the GEM network in unseen environments significantly.
+For example, in the zero-shot sim-to-real setting of real-robot experiments, the average success rate over different skills is improved by 22%.
+According to the results of GEM-O and GEM-S, gradually removing the color perturbation achieves better performance than constantly applying strong color perturbation.
+This is because the visual appearance information is useful in scenarios like distinguishing two closely placed objects.
+For example, in the Insert task, the network needs to discriminate the inserted object and container, and solely relying on object geometry is difficult to meet this requirement.
+Therefore, we can find from Table III that the Pick success rates of GEM-S are significantly weaker than GEM-O.
+TABLE IV: Comparison between point and image observations.
 […]
-0.05 0.92 0.96 0.79 0.74
-0.10 0.90 0.97 0.80 0.71
-0.25 0.81 0.92 0.73 0.48
-0.50 0.00 0.00 0.00 0.00
-V-D3 Analysis on Visual Servo tracking
-In GEM, we utilize the visual servo tracking module to convert the dynamic object manipulation problem into static object manipulation in the view of the recurrent probabilistic network.
-This study analyzes the tracking stability and velocity upper limit in real-world manipulation.
-Specifically, we adjust the conveyor belt to various speeds and make the robot hand track a moving object based on the visual servo tracking module.
-It is observed that the maximum stable tracking speed in simulation is approximately 0.25 m/s, and it reaches about 0.3 m/s in the real world.
+Image 0.71 0.35 0.41 0.39
+Point 0.93 0.89 0.95 0.87
+V-C2 Analysis on Point based Representation
+According to Table III, the GEM policy can generalize from simulated data to real-world manipulation to some extent even without the appearance noise annealing strategy.
+This characteristic is attributed to the 3D point observation of GEM rather than using images.
+As analyzed in Section III-D, 3D points directly deliver geometry, which is inherently of higher predictivity for estimating manipulation actions.
+Therefore, by utilizing 3D point observation, GEM naturally prioritizes geometry information over visual appearance, and our developed appearance noise annealing strategy strengthens this prioritization.
+By contrast, if taking images as input, the network has to infer 3D geometry from visual appearance, meaning that the network must rely on exploiting visual appearance features.
+In this experiment, we study the performance gain brought by adopting point observation compared with employing image observation.
+When the network input is of the image modality, the GEM network encoders are switched to DINOv2 [36], which is a popular image backbone with discriminative representation.
+The experimental results of adopting point and image observations are presented in Table IV.
+According to the results, we can find that the models taking 3D points as input outperform the ones with images significantly.
+For example, in the zero-shot sim-to-real setting, the point-based policy surpasses the image-based policy by 48%, which confirms the importance of using 3D point observation.
+Figure 11: This figure visualizes some instance segmentation cases during manipulation. We render the segmented points in orange and highlight the segmentation mask with an additional orange box.
+V-C3 Analysis on Object Segmentation
+As depicted in Fig.
+6, the GEM network predicts the 3D point segmentation mask of the target object to manipulate, and this segmentation mask is employed to generate tracking actions.
+Therefore, whether the network can produce a coarse mask of the target object is important for stably tracking the target object.
+To enable this stable tracking, we need promising segmentation mask annotations in training data.
+Thankfully, since we employ simulators to produce demonstrations, we can obtain precise segmentation labels automatically.
+Besides, since the GEM network primarily utilizes geometry feature to identify the target object, we find the network trained solely using simulation data generalizes to the real world well.
+To validate this, we collect a real-world segmentation dataset consisting of 1800 point scenes based on the assets shown in Fig.
+9.
+Then, we utilize SAM2 [40], a state-of-the-art open-world segmentation model, to generate segmentation labels efficiently.
+The obtained labels are employed to validate the GEM network.
+Following the testing protocol of the COCO benchmark [27], the GEM network achieves a Mask AP of 86.5%, which meets the requirement of stable tracking.
+Moreover, we visualize several target object segmentation examples in Fig.
+11, where the segmentation mask is highlighted in orange.
+The results demonstrate promising segmentation performance.
+As shown in Fig.
+11 (a), in the Pick or Rotate task that the operation involves only a single object, the network accurately segments that object.
 In Fig.
-17, we visualize the dynamic curves of the control target signal and corresponding end-effector position.
-We can find that the end-effector approaches the moving object fast and tracks it stably under the velocity of 0.2 m/s.
-However, the end-effector fails to track the object when the velocity is 0.6 m/s due to its speed limit.
-Afterwards, we validate the real-world success rates of GEM in the settings of different moving speeds.
-The results are reported in Table IV.
-Comparing the $1_{\rm st}$ and $2_{\rm nd}$ rows of results, we can observe that the success rates are similar when the conveyor belt runs at a low speed.
-When the speed is 0.25 m/s, the success rates drop due to stricter manipulation precision requirements.
-When the speed arrives at 0.50 m/s, the robot hand cannot perform any manipulation task because it cannot keep pace with the movement of the target object.
-TABLE V: Analysis on the effect of the memory encoder.
+11 (b), for the Put or Insert task that a gripper places an item into a container, the network successfully identifies the item based on its geometry characteristics.
+TABLE V: Analysis on the effect of the memory module.
 […]
-$\times$ 0.72 0.60 0.69 0.55
-$\checkmark$ 0.87 0.78 0.85 0.74
-V-D4 Analysis on Memory Modeling
-As discussed in Section III-D, enabling a network to access historical memory is important for reducing action prediction ambiguity.
-To this end, we devise the memory encoder shown in Fig.
-8.
-In experiments, we find that without this encoder, the robot arm could oscillate at the initial position of manipulation without approaching the target, leading to task failure.
+$\times$ 0.86 0.76 0.90 0.82
+$\checkmark$ 0.93 0.89 0.95 0.87
+V-C4 Analysis on Memory Modeling
+In dynamic object manipulation, modeling historical memory enhances the policy’s awareness of both the object’s motion dynamics and task progress.
+Therefore, we devise the memory module shown in Fig.
+6.
+In experiments, we find that without this module, the robot arm could oscillate at the initial position of manipulation without approaching the target, leading to task failure.
 This phenomenon is more frequent when a single network learns multiple skills.
 This is because the policy fails to distinguish whether the manipulation operation has not started or has been completed.
-Our designed memory encoder can effectively address this problem by enabling the network to know which actions have been performed.
-We compare the performances of GEM with and without this memory encoder and report the results in Table V.
-According to the results, we can observe that the memory encoder significantly boosts success rates.
-To more thoroughly analyze how long a sequence the memory encoder can memorize, we design a toy experiment of reciting a random digit sequence.
-Specifically, we update the structure of this memory encoder to support accepting a single digit as input and predicting the next digit per step.
-Then, we randomly generate a sequence of $n$ digits and train the memory encoder to recite this sequence.
+Our designed memory module can effectively address this problem by enabling the network to know which actions have been performed.
+We compare the performances of GEM with and without this memory module and report the results in Table V.
+According to the results, we can observe that the memory module significantly boosts success rates.
+To more thoroughly analyze how long a sequence the memory module can memorize, we design a toy experiment of reciting a random digit sequence.
+Specifically, we update the structure of this memory module to support accepting a single digit as input and predicting the next digit per step.
+Then, we randomly generate a sequence of $n$ digits and train the memory module to recite this sequence.
 After training, we test how many digits can be correctly recited.
-After repeating this experiment by setting different values of $n$, we find that the memory encoder can correctly recite a very long sequence like 1000 digits after sufficient epochs of training.
+After repeating this experiment by setting different values of $n$, we find that the memory module can correctly recite a very long sequence like 1000 digits after sufficient epochs of training.
 Such a memory capability is strong enough for existing embodied policies, because current manipulation tasks usually involve at most tens of times of policy inference.
@@
 […]
-Direct 0.81 0.73 0.79 0.67
-Probabilistic 0.87 0.78 0.85 0.74
-V-D5 Analysis on Network Output
-As explained in Section III-E, the output of the recurrent probabilistic network is a Gaussian mixture distribution to represent the predicted action probabilistic distribution.
+Direct 0.88 0.80 0.91 0.79
+DP 0.81 0.76 0.85 0.63
+FM 0.86 0.77 0.88 0.75
+Probabilistic 0.93 0.89 0.95 0.87
+V-C5 Analysis on Network Output
+As explained in Section IV-D, the output of the GEM network is a GMM-based probabilistic distribution to capture the multi-modal nature of actions.
 This is different from previous works that directly regress the actions to execute [13, 7].
-This part analyzes the benefits of our probabilistic output.
-We first compare the performances of networks based on direct action prediction and probabilistic action prediction.
+This part analyzes the benefits of adopting this GMM-based probabilistic output.
+We first compare the performances of the GEM network outputs based on direct action prediction, diffusion process (DP), flow matching (FM), and GMM.
 The results are reported in Table VI.
-We can observe that the recurrent probabilistic network based on probabilistic action prediction outperforms the one using direct action prediction, because the multimodal ambiguity problem is alleviated.
-Notably, we argue that the results in Table VI do not sufficiently reveal the value of probabilistic action prediction.
-This is because the recurrent probabilistic network only observes a single target object due to the object-centric geometry perception block, and the actions in the collected demonstrations are always to manipulate this single object.
-There is no case like the one shown in Fig.
-4, where multiple target objects appear simultaneously in the network observation.
-To bridge this gap, we develop a simulated environment as shown in Fig.
-18.
+We can observe that the GMM-based network outperforms the other counterparts employing the compared output strategies, because the GMM output effectively alleviates the multimodal ambiguity problem.
+In addition, it is found that the performances of the DP and FM based outputs are inferior to direct action prediction, though DP and FM can also handle the multimodal ambiguity problem.
+After analysis, we infer this is because the denoising process in DP and FM is challenging for optimization, so a large training batch size is demanded to achieve promising manipulation action generation.
+In this experiment, since the training demonstration volume is limited, we cannot set such a large batch size.
+Thus, the obtained success rates are limited.
+According to these findings, we can summarize that the requirement of a large training batch size makes DP and FM less flexible than our developed GMM-based network output.
+Notably, the results in Table VI do not fully demonstrate the advantage of probabilistic action prediction.
+This is because the objects to manipulate arrive sequentially in dynamic object manipulation on conveyor belts.
+Hence, at a single moment, the policy usually observes only one object that is suitable for manipulation.
+As a result, the multi-modal ambiguity that the probabilistic head is designed to resolve remains limited.
+To more sufficiently reveal the value of probabilistic action prediction, we develop a simulated environment as shown in Fig.
+12.
 We use the teleoperation handle shown in Fig.
 8 (b) to control the robot hand to collect 50 demonstrations, where we randomly pick up one of these two blocks and put it into the green container.
-To study multi-target ambiguity, we remove the object-centric geometry perception block in GEM and make the recurrent probabilistic network take images as input.
-The feature extractor is changed to DINOv2.
 After training using the collected demonstrations, we compare the success rates of the networks based on direct action prediction and probabilistic action prediction.
@@
 […]
-$\times$ 0.35 0.28 0.20 0.16
-$\checkmark$ 0.87 0.78 0.85 0.74
-V-D6 Analysis on Hybrid Action Control
-The hybrid action control block combines visual servo control and learnable network control to conduct dynamic object manipulation.
-This experiment studies how this strategy contributes to success rates compared with the common design, i.e., solely based on learnable network control.
-The performance comparison results between GEMs with and without using the visual servo control are presented in Table VII.
-We can find that the visual servo control, which converts the original dynamic object manipulation problem into static manipulation, is critical for the overall task completion success rates.
-TABLE VIII: Analysis on the effect of simulated demonstrations.
+$\times$ 0.37 0.31 0.45 0.30
+$\checkmark$ 0.93 0.89 0.95 0.87
+V-C6 Analysis on Manipulation Action Decomposition
+In GEM, we design a strategy that decomposes the manipulation actions into interaction actions and tracking actions, where the interaction actions are decoded from the predicted GMM distribution and the tracking actions are produced based on the point segmentation result.
+Within this strategy, the tracking actions control the robot end-effector to follow the motion of the target object.
+In this way, the pose of the end-effector keeps stationary relative to the target object, and thus the interaction actions can be regarded as operating a static object.
+This design significantly simplifies this task.
+The experimental results in Table I and Table II indicate the effectiveness of this strategy.
+In these two tables, the compared policies marked with an asterisk (*) employ the action decomposition strategy, whereas the ones without an asterisk do not.
+We can find that the developed strategy improves their success rates by large margins.
+In addition, we also present the results of GEM without and with the manipulation action decomposition strategy in Table VII.
+According to the results, it is observed that this strategy is important for the overall performance of GEM.
+Figure 13: This figure visualizes the tracking dynamics of the end-effector under different target velocities.
+We can find that the end-effector stably tracks objects moving at 0.2 m/s but fails to keep up with objects moving at 0.6 m/s.
+TABLE VIII: Manipulation performance in various moving speeds.
 […]
-$\times$ 0.28 0.45 0.15 0.08
-$\checkmark$ 0.92 0.96 0.79 0.74
-V-D7 Analysis on Data Efficiency
-As explained in Section IV-F, since the recurrent probabilistic network takes uncolored 3D points as input, the observations in simulators become similar to real-world observations.
-Therefore, the simulated data can directly benefit real-world performances.
-Correspondingly, we collect 6,600 demonstrations in our built simulation environment.
-The policy trained using these simulated demonstrations should significantly favor the real-world results.
-To verify this issue, we compare the real-world performances of the GEM policies with and without being pre-trained with these simulated demonstrations.
-Both the two policies are tuned using the 80 real-world demonstrations described in Section V-B before the real world evaluation.
-The experimental results are presented in Table VIII.
-We can find that the simulated data boosts the success rates of various skills by large margins.
-According to the first row of results, only utilizing the 80 real-world demonstrations is insufficient to derive a promising policy, and collecting more real-world demonstrations is labor-intensive.
-Thankfully, simulated demonstrations can tackle this data scarcity problem because simulated data generation can be efficient.
-V-E Generalization Analysis
-The various designs in GEM are to improve the generalization of this manipulation system with minimal cost.
-This subsection conducts experiments to validate the generalization capabilities of GEM from different perspectives.
-Figure 19: To study the generalization performance of GEM across different environment backgrounds, we randomly augment the appearance of the conveyor belt, and these generated appearances are unseen in demonstrations.
+0.05 0.92 1.00 0.91 0.87
+0.10 0.94 1.00 0.90 0.86
+0.25 0.86 0.98 0.85 0.63
+0.50 0.00 0.00 0.00 0.00
+V-C7 Analysis on Object Tracking
+This work studies dynamic object manipulation on conveyor systems, where the target objects could move at various velocities.
+As discussed in the above paragraphs, we design a strategy that decomposes the manipulation actions into interaction actions and tracking actions to handle the diverse motion dynamics of target objects.
+In this part, we validate the performance of GEM on manipulating objects of different velocities.
+First of all, we analyze the tracking stability and velocity upper limit in real-world manipulation.
+Specifically, we adjust the conveyor belt to various speeds and make the robot hand track a moving object based on the visual servo tracking module.
+It is observed that the maximum stable tracking speed in simulation is approximately 0.25 m/s, and it reaches about 0.3 m/s in the real world.
+In Fig.
+13, we visualize the dynamic curves of the control target signal and corresponding end-effector position.
+We can find that the end-effector approaches the moving object fast and tracks it stably under the velocity of 0.2 m/s.
+However, the end-effector fails to track the object when the velocity is 0.6 m/s due to its speed limit.
+Subsequently, we validate the success rates of GEM on manipulating objects of different moving speeds in the simulated in-domain experiment setting.
+The results are reported in Table VIII.
+Comparing the $1_{\rm st}$ and $2_{\rm nd}$ rows of results, we can observe that the success rates are similar when the conveyor belt runs at a low speed.
+When the speed is 0.25 m/s, the success rates drop because the robot velocity limit prevents the end-effector from maintaining stable tracking while simultaneously performing precise manipulation.
+When the speed arrives at 0.50 m/s, the robot hand cannot perform any manipulation task because it cannot keep pace with the movement of the target object.
+V-D Generalization Analysis
+Thanks to the many designs in GEM, it shows strong generalization in various unseen scenarios, making it suitable for real applications.
+This subsection analyzes the generalization capabilities of GEM from different perspectives.
+Figure 14: To study the background generalization of GEM while excluding the influence of irrelevant factors, we randomly augment the appearance of the conveyor belt in Isaac Gym.
+The augmented appearances are unseen in demonstrations.
 Two augmented samples are visualized in this figure.
@@
 […]
-Seen 0.88 0.98 0.83 0.80
-Unseen 0.91 0.96 0.83 0.81
-V-E1 Generalization across Environment Backgrounds
-Thanks to the rapid progress of instance segmentation [57, 66], obtaining a segmentation model meeting manipulation precision requirements has become highly cost-effective.
-Therefore, we design the object-centric geometry perception block in GEM that produces uncolored 3D points appended with instance IDs as the input to the recurrent probabilistic network.
-In this way, regardless of changes in the environment background, the input to the network remains similar.
-Therefore, our method should present strong generalization across different environmental backgrounds.
-To validate this characteristic, we test the manipulation performance of GEM in different unseen backgrounds by changing the appearance of the conveyor belt.
-Since changing conveyor belts is expensive to realize in the real world, we conduct this experiment in simulation.
-As shown in Fig.
-19, we randomly augment the colors of the conveyor belt.
-Since all demonstrations are collected using the conveyor belt illustrated in Fig.
-9 (a), these augmented conveyor belts are unseen in demonstrations.
-The manipulation success rates of GEM in the original seen environment and the augmented unseen environments are presented in Table.
-IX.
-We can observe that the performances of GEM are similar in seen and unseen environments, which is consistent with the expectation.
+Seen 0.94 1.00 0.90 0.86
+Unseen 0.95 1.00 0.88 0.85
+V-D1 Generalization across Environment Backgrounds
+In applications, various scenarios show different environment backgrounds.
+It is expensive to collect demonstrations in every scenario to tune a policy.
+Thus, we expect that a policy can generalize to unseen scenarios for supporting more flexible deployment.
+Considering this expectation, we develop GEM, which prioritizes geometry structure and decreases the importance of visual appearance.
+In this part, we analyze the generalization capability of GEM in unseen backgrounds.
+Although we develop the out-of-domain simulation testing protocol, where the policy is trained using demonstrations from Isaac Gym and tested in Isaac Sim, this protocol involves extra influence factors such as simulation parameters.
+To exclude irrelevant factors, we choose to randomly augment the appearance of the conveyor belt in Isaac Gym, as visualized in Fig.
+14.
+In this way, the policy is trained in Isaac Gym and validated in Isaac Gym, where the environment background is the only difference.
+We evaluate the performance of GEM and present the results in Table IX.
+According to the results, the performances of GEM are similar in environments with seen and unseen backgrounds, which is consistent with the expectation.
 In addition, the subsequent sim-to-real experiment in Table XIII indicates that the GEM trained using only simulated data can generalize to real-world manipulation despite the appearance gap between the simulated and real-world environments.
@@
 […]
-Simulated Linear 0.88 0.98 0.83 0.80
-Simulated S-curve 0.91 0.97 0.80 0.77
-Real Linear 0.92 0.96 0.79 0.74
-Real Random curve 0.88 0.98 0.75 0.66
-V-E2 Generalization across Motion Patterns
-As the visual servo control in the hybrid action control block transforms the original dynamic object manipulation problem into a static manipulation problem, the GEM policy should can generalize to manipulating objects in diverse motion patterns, e.g., various moving velocities and trajectories.
-The experimental results in Table IV have confirmed that the manipulation success rates remain similar when the conveyor belt speed is adjusted but does not surpass the robot hand tracking velocity limit.
+Simulated Linear 0.94 1.00 0.90 0.86
+Simulated S-curve 0.95 1.00 0.88 0.81
+Real Linear 0.98 1.00 0.93 0.89
+Real Random curve 0.97 1.00 0.90 0.75
+V-D2 Generalization across Motion Patterns
+As the tracking actions generated based on the manipulation action decomposition strategy transform the original dynamic object manipulation problem into a static object manipulation problem, the GEM policy should be able to generalize to manipulating objects in diverse motion patterns, e.g., various moving velocities and trajectories.
+The experimental results in Table VIII have confirmed that the manipulation success rates remain similar when the conveyor belt speed is adjusted but does not surpass the robot hand tracking velocity limit.
 In this experiment, we further study this problem by altering the object moving trajectories.
@@
 We report the experiment results in Table X.
-Comparing the performance between linear and curved object movement trajectories, we can conclude that making objects move along an unknown curved trajectory does not harm the manipulation success rates significantly.
+Comparing the performance between linear and curved object movement trajectories, we can find that making objects move along an unknown curved trajectory does not harm the manipulation success rates significantly, especially for the tasks that do not require precise operation like Put.
 Therefore, GEM is robust to different motion patterns of moving objects.
 Figure 16: This figure visualizes the four kinds of robot hands to conduct cross-embodiment evaluation in the simulated and real environments.
-TABLE XI: Generalization study on object geometries.
+TABLE XI: Generalization study on unseen objects.
 […]
-Seen 0.92 0.96 0.79 0.74 0.85
-Unseen 0.80 0.90 0.67 0.59 0.74
-V-E3 Generalization across Object Geometries
-By collecting demonstrations based on objects in diverse geometries and training the recurrent probabilistic network using these demonstrations, GEM can generalize to different geometries.
-The experiment results in Table I suggest that GEM can perform various manipulation skills on unseen objects with promising success rates.
-Table II reports the performance of GEM on manipulating seen objects in real-world tasks, where the results are similar to the ones in simulated environments.
-In this experiment, we further reveal the real-world success rates on manipulating the right nine unseen objects in Fig.
-11.
-The recurrent probabilistic network validated in this experiment is the same as the one in Table II, which is first pre-trained using the 6,600 simulated demonstrations and then fine-tuned with 20 real-world demonstrations per skill.
+Seen 0.94 1.00 0.90 0.86 0.93
+Unseen 0.89 0.98 0.87 0.85 0.90
+V-D3 Generalization across Unseen Objects
+Besides environment backgrounds and object motion patterns, generalizing to manipulating unseen objects is also important in real application scenarios.
+The experimental results reported in Table I and Table II have suggested that the GEM policy can operate novel objects.
+This advantage is thanks to the training demonstrations collected on objects of diverse geometries, enabling the GEM network to generalize to out-of-domain object geometries.
+However, there exist extra influence factors in these two experiments, such as inconsistent simulation parameters and new robot embodiments.
+In this experiment, we further analyze the performance of GEM on manipulating unseen objects after excluding irrelevant factors.
+To this end, we import the novel 3D assets shown in Fig.
+10 (b) into Isaac Gym to evaluate the GEM trained based on the demonstrations from Isaac Gym.
 The experiment results are reported in Table XI.
-We can find that although the network is only tuned with 20 demonstrations per skill, it still generalizes well to unseen objects in diverse geometries.
-In Section V-F, we show that if we fine-tune the network using 100 Pick demonstrations, it achieves a success rate of over 97% in an unseen application environment.
-V-E4 Generalization across Robot Embodiments
-Thanks to the visual servo control, the end-effector moves to a fixed track position relative to the target object before that the recurrent probabilistic network begins generating manipulation actions.
-Additionally, the robot hand is controlled by the predicted end-effector relative offset based on inverse kinematics.
-Therefore, GEM is not restricted within a single robot embodiment.
+We can observe that the success rates of GEM in manipulating out-of-domain objects are similar to the ones of operating in-domain objects.
+This is because GEM primarily exploits geometry information, and the basic 3D geometries of objects are limited despite their diverse visual appearances.
+The 3D assets depicted in Fig.
+7 (a) have covered many common basic geometries.
+V-D4 Generalization across Robot Embodiments
+Because of the developed manipulation action decomposition strategy, the end-effector moves to a preset position relative to the target object before the predicted interaction actions are applied.
+Additionally, the manipulation actions control the joints of a robot based on inverse kinematics.
+Therefore, GEM is not restricted to a single robot embodiment.
 To evaluate the robot embodiment generalization, we conduct experiments using four kinds of robot hands in simulated and real-world environments, as shown in Fig.
-21.
-In the simulation, the policy trained using the simulated demonstrations collected from Franka Panda is directly applied to the UR10e robot hand.
-In the real-world experiment, the policy tuned based on the demonstrations collected with Piper is deployed on the ARX5 robot hand to conduct manipulation.
+16.
+In the simulation based on Isaac Gym, the policy trained using the simulated demonstrations collected from Franka Panda is directly applied to the UR10e robot hand.
+In the real-robot experiment, the policy tuned based on the demonstrations collected with Piper is deployed on the ARX5 robot hand to conduct manipulation.
 The experiment results are reported in Table XII.
-It is observed that employing a robot hand to collect demonstrations and adopting another robot hand to conduct manipulation evaluation achieves similar performance compared with the setting where the demonstration collection and validation are based on the same robot hand.
+It is observed that employing a robot to collect demonstrations and adopting another robot to conduct manipulation evaluation achieves similar performance compared with the setting where the demonstration collection and validation are based on the same robot.
 Hence, the robot embodiment generalization capability of GEM is confirmed.
@@
 […]
-Franka Franka 0.88 0.98 0.83 0.80
-Franka UR10e 0.91 0.97 0.85 0.79
-Piper Piper 0.92 0.96 0.79 0.74
-Piper ARX5 0.90 0.95 0.77 0.69
-TABLE XIII: Analysis on sim-to-real generalization.
+Franka Franka 0.94 1.00 0.90 0.86
+Franka UR10e 0.95 1.00 0.88 0.86
+Piper Piper 0.98 1.00 0.93 0.89
+Piper ARX5 0.96 1.00 0.94 0.85
+TABLE XIII: Sim-to-real performance of GEM.
 […]
-$\times$ 0.76 0.83 0.55 0.48
-$\checkmark$ 0.92 0.96 0.79 0.74
-V-E5 Sim-to-Real Generalization
-The input to the recurrent probabilistic network is uncolored 3D points, which do not exhibit significant differences between simulated and real environments.
-Therefore, GEM solely based on simulated demonstrations can generalize to the real world to perform operations.
-To validate this characteristic, we report the real-world manipulation success rates of GEMs with and without tuning in Table XIII.
-According to the results, we can observe that the GEM trained using only simulated data can manipulate objects in the real world, though its success rates are lower than the GEM tuned with 20 demonstrations per skill.
-Through analysis, we find this is caused by two reasons.
-First of all, without tuning, the test objects are all unseen in demonstrations, which increases the manipulation challenge.
-Secondly, in a robot control system, there is a control execution delay between the input signal and the system response.
-This control execution delays are different between simulated and real robots, further enhancing the difficulty.
-V-F Real Application Deployment
+Real_tune 0.98 1.00 0.93 0.89 0.95
+Real_zero 0.91 0.98 0.80 0.79 0.87
+V-D5 Sim-to-Real Generalization
+According to the aforementioned experiments, GEM can generalize to novel environment backgrounds, motion patterns, target objects, and robot embodiments.
+These characteristics enable the possibility of generalizing to real-world tasks based on simulated demonstrations.
+To validate this issue, we design two real-robot experiment settings, i.e., tuned sim-to-real and zero-shot sim-to-real.
+The experimental results of the tuned sim-to-real setting are presented in Table II.
+The results suggest that GEM achieves significantly superior sim-to-real generalization with a handful of tuning demonstrations.
+The performance of GEM in the zero-shot sim-to-real setting is reported in Table XIII.
+According to the results, it is observed that the GEM policy trained using only simulated data can manipulate objects in the real world, and its success rates are still promising.
+This strong generalization is attributed to the design that GEM prioritizes geometry information and decreases the importance of visual appearance.
+We also analyze the reasons of the slight performance deterioration.
+An important factor is there exists a control execution delay between the input signal and the robot system response.
+This control execution delays are different between simulated and real robots, causing a sim-to-real execution gap.
+Tuning the policy with a few real-robot demonstrations can bridge this execution delay gap.
+Therefore, the performance of GEM in the tuned sim-to-real setting is stronger.
+V-E Real Application Deployment
 Different from lab environments, it is often infeasible to collect demonstrations in real application scenes because of the uninterrupted normal workflows and strict confidentiality requirements.
@@
 Since onsite data collection is infeasible in this canteen, we teleoperate 100 demonstrations using the manipulation platform shown in Fig.
-5.
-Then, we utilize these 100 demonstrations to fine-tune the GEM policy that has been pre-trained with the 6,600 simulated data.
+8.
+Then, we utilize these 100 demonstrations to fine-tune the GEM policy that has been pre-trained with the simulation demonstrations.
 Finally, we deploy the fine-tuned GEM in this canteen to conduct spoon collection.
 We evaluate GEM over a seven-day period, processing more than ten thousand spoons, and measure a final manipulation success rate of 97.2%.
-This marks the first real application of imitation learning in dynamic object manipulation.
 VI Conclusion
-This paper has first developed an entropy-based theoretical framework to analyze the data requirements of imitation learning.
-Based on this framework, we have proposed GEM that reduces observation complexity and action prediction ambiguity to achieve strong generalization.
-Through extensive experiments, GEM has successfully generalized across diverse environment backgrounds, motion patterns, object geometries, and robot embodiments, achieving a success rate of over 97% in a real-world application without using any in-scene demonstration.
-These results have demonstrated that our theoretical insights effectively guide the design of an efficient manipulation system, offering a practical solution for industrial automation with reduced data costs.
+This paper has first pointed out that embodied action generation concerns more about the 3D geometry structure of environments than visual appearance.
+Based on this insight, we have proposed the policy GEM for dynamic object manipulation on conveyor systems.
+The key design principle of this policy is prioritizing the exploitation of 3D geometry information and decreasing the reliance on 2D visual information.
+In addition, we have developed simulation platforms of the explored tasks to generate training demonstrations efficiently.
+Since the simulators can simulate the geometry-based interaction dynamics among objects well and GEM primarily uses geometry information, GEM achieves promising sim-to-real generalization.
+We have conducted extensive experiments to analyze the characteristics of GEM.
+The results indicate that GEM has successfully generalized across diverse environment backgrounds, motion patterns, object geometries, and robot embodiments, achieving a success rate of over 97% in a real-world application without using any in-scene demonstration.
+These results have demonstrated that our policy design principle effectively guides the design of an efficient manipulation system, offering a practical solution for industrial automation.